### v1

# The Edge-of-Reach Problem in Offline Model-Based Reinforcement Learning

Abstract
Offline reinforcement learning aims to enable agents to be trained from pre-collected datasets, however, this comes with the added challenge of estimating the value of behavior not covered in the dataset. Model-based methods offer a solution by allowing agents to collect additional synthetic data via rollouts in a learned dynamics model. The prevailing theoretical understanding is that this can then be viewed as online reinforcement learning in an approximate dynamics model, and any remaining gap is therefore assumed to be due to the imperfect dynamics model. Surprisingly, however, we find that if the learned dynamics model is replaced by the true error-free dynamics, existing model-based methods completely fail. This reveals a major misconception. Our subsequent investigation finds that the general procedure used in model-based algorithms results in the existence of a set of edge-of-reach states which trigger pathological value overestimation and collapse in Bellman-based algorithms. We term this the edge-of-reach problem. Based on this, we fill some gaps in existing theory and also explain how prior model-based methods are inadvertently addressing the true underlying edge-of-reach problem. Finally, we propose Reach-Aware Value Learning (RAVL), a simple and robust method that directly addresses the edge-of-reach problem and achieves strong performance across both proprioceptive and pixel-based benchmarks.
Offline Reinforcement Learning, Model-Based Reinforcement Learning
1 Introduction
Figure 1: Boundaries of the states reachable at each timestep in the $k$-step rollouts from $\mathcal{D}_{\textrm{offline}}$ as used in offline model-based RL. In this paper we investigate a hitherto unrecognized pathology in offline model-based methods whereby edge-of-reach states - those only reachable at the final step of model rollouts (in red) - can trigger value overestimation and collapse of training.
Standard online reinforcement learning (RL) requires the collection of large amounts of on-policy data throughout training. This can be both costly and unsafe, and hence represents a barrier against applying RL in domains such as healthcare (Tang & Wiens, 2021; Shiranthika et al., 2022) or robotics (Kumar et al., 2021; Chebotar et al., 2021; Ball et al., 2021), and also against scaling RL to more complex problems. Offline RL (Levine et al., 2020; Ernst et al., 2005) aims to remove this need for online data collection by enabling agents to be trained on pre-collected datasets. The use of large pre-existing datasets has driven significant recent advances in supervised learning (Brown et al., 2020), and one hope is that offline RL may provide a route for similar scaling in RL (Kumar et al., 2023).Code open-sourced at: github.com/anyasims/edge-of-reach.
The central challenge in offline RL is dealing with value estimation for counterfactual actions not present in the dataset, termed the out-of-sample action problem (Kostrikov et al., 2022). [^1] Also referred to as the out-of-distribution action (Kumar et al., 2019) or action distribution shift problem (Kumar et al., 2020). A naïve approach results in inaccurate values at out-of-sample state-actions being propagated through the entire state-action space, leading to pathological training dynamics often exhibited as extreme value overestimation (Kumar et al., 2019). There have been many proposals to resolve this, with methods largely falling into one of two categories: model-free (Kostrikov et al., 2022; Kumar et al., 2020; Fujimoto et al., 2019; Kumar et al., 2019; An et al., 2021; Fujimoto & Gu, 2021) or model-based (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023; Lu et al., 2022).
Model-free methods typically address this issue by applying a form of conservatism or constraint to avoid using out-of-sample actions in the Bellman update. In contrast, model-based methods, which are the focus of this paper, generally aim to resolve the issue by using the offline dataset to fit an approximate dynamics model (Sutton, 1991; Janner et al., 2019), and then allowing the agent to collect additional data in this learned dynamics model through fixed horizon rollouts from the original offline dataset (the general MBPO procedure (Janner et al., 2019), see Algorithm 1). The current understanding is that this converts the offline learning problem into online RL in an approximate dynamics model, thus solving the out-of-sample problem by allowing the agent to collect corrective feedback for any misestimated values. With this understanding, model-based methods typically attribute any remaining difficulties to errors in the learned dynamics model, and algorithms are subsequently motivated by avoiding exploitation of model errors using some notion of dynamics uncertainty.
This understanding of model-based approaches naturally implies that reducing dynamics model errors should improve performance. However, we begin with a surprising demonstration that existing offline model-based methods completely fail when trained with the true error-free dynamics model (see Table 1 below). This exposes a significant misconception in the current understanding of model-based offline RL and prompts us to reevaluate our understanding of how prior methods work. On investigation, we find that the use of short rollouts in these algorithms (as opposed to full horizon trajectories as collected in online RL) causes a critical and previously unnoticed problem. More specifically, it leads to the existence of states which, under any policy, can only be reached in the final rollout step (see red in Figure 1). These edge-of-reach states appear in the training distribution as nextstates, and are hence used in Bellman updates, but even with the ability to collect unlimited additional data, the agent is never able to observe the outcome of actions from them. The result is pathological overestimation of values, similar to that seen in model-free methods. We term this the edge-of-reach problem. Thus, contrary to common understanding, the out-of-sample problem central in model-free methods is not resolved by a model-based approach.
This edge-of-reach problem reveals strong connections between the two previously disjoint subfields of offline model-free and model-based RL. With this new understanding, we propose Reach-Aware Value Learning (RAVL), a robust scalable solution that takes an existing solution to the central out-of-sample problem in model-free literature, and directly applies it to the edge-of-reach problem in model-based methods (both proprioceptive and pixel-based). We first analyze RAVL on simple toy environments, and then show that it scales to the standard D4RL (Fu et al., 2020) and the V-D4RL (Lu et al., 2023) benchmarks. Furthermore, we anticipate that this unified understanding of model-based and model-free algorithms may pave the way for porting further strengths of existing model-free algorithms into model-based approaches.
Code is open-sourced at: github.com/anyasims/edge-of-reach
To summarize, our contributions are as follows:
  • We expose a significant misconception in the current understanding of offline model-based RL (Section 4.1). This leads to our hypothesis and definition of the edge-of-reach problem (Section 4.2). We provide evidence that this is the true underlying source of issues on the standard D4RL (Fu et al., 2020) benchmark (Section 4.4), before then constructing a simple environment which allows us to demonstrate and verify this hypothesis (Section 5).
  • Based on this, we propose Reach-Aware Value Learning (RAVL), a natural solution to the edge-of-reach problem (Section 6). We extensively evaluate RAVL across the D4RL and pixel-based V-D4RL (Lu et al., 2023) benchmarks (Section 7), including careful analysis into what RAVL is learning, demonstrating that it directly targets the edge-of-reach problem as intended (Section 7.1) and achieves state-of-the-art performance.
  • Finally, we give a more accurate explanation of the efficacy of prior methods, showing they unintentionally address the edge-of-reach problem (Section 7.4). [^2] Throughout the paper, please refer to Table 9 for the naming conventions used in experiments.
2 Background
Reinforcement Learning. We consider the standard RL framework (Sutton & Barto, 2018), in which the environment is formulated as a Markov Decision Process, $M=(\mathcal{S},\mathcal{A},T,R,\mu_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, $T(s^{\prime}|s,a)$ and $R(s,a)$ denote the transition and reward dynamics, $\mu_{0}$ the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. The goal in reinforcement learning is to learn a policy $\pi(a|s)$ that maximizes the expected discounted return $\mathbb{E}_{\mu_{0},\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$.
Actor-Critic Algorithms. The broad class of algorithms we consider are actor-critic (Konda & Tsitsiklis, 1999) methods which jointly optimize a policy $\pi$ and state-action value function ($Q$-function). Actor-critic algorithms iterate between fitting the $Q$-function to the current policy (policy evaluation) and improving the policy relative to the $Q$-function (policy improvement). The policy evaluation step relies on approximate dynamic programming with Bellman updates, updating the $Q$-function using some dataset $\mathcal{D}$ as follows:
[…]
Offline RL and the Out-of-Sample Problem. In online RL, the agent is able to collect new on-policy data throughout training. By contrast, in the offline setting, the algorithm has access only to a fixed offline dataset consisting of transition tuples $\mathcal{D}_{\textrm{offline}}=\{(s,a,r,s^{\prime})^{(i)}\}_{i=1,\dots,N}$ collected by some behavioral policies $\pi^{\beta}$. The central problem in offline learning is the out-of-sample [^3] We use the terms ‘out-of-sample’ and ‘out-of-distribution’ interchangeably, as is done in the literature. action problem: Target actions being sampled on-policy $a^{\prime}\sim\pi^{j}$ means state-actions $(s^{\prime},a^{\prime})$ may not appear as $(s,a)$ in the dataset (see Equation 1). Values at these out-of-sample state-actions are therefore never themselves updated and hence likely to be arbitrarily misestimated, but yet they are used for the targets in Bellman updates. Updates therefore propagate this misestimation to previous (in-sample) state-actions leading to a complete breakdown in $Q$-learning. Further, the $\max$ operation implicit in the policy improvement serves to pick out the highest misestimated $Q$-values such that $Q$-values become extremely overestimated, often increasing unboundedly over training (Levine et al., 2020; Kumar et al., 2019).
Offline Model-Based RL Model-based methods (Sutton, 1991) aim to solve the out-of-sample issue by allowing the agent to collect additional synthetic data in a learned dynamics model $\widehat{M}=(\mathcal{S},\mathcal{A},\widehat{T},\widehat{R},\mu_{0},\gamma)$. $\widehat{T}(s^{\prime}|s,a)$ and $\widehat{R}(s,a)$ denote the approximate transition and reward functions, and are commonly realized as a deep ensemble (Chua et al., 2018; Lakshminarayanan et al., 2017). Methods typically use the MBPO procedure (Janner et al., 2019) shown in Algorithm 1, with synthetic data sampled as short $k$-step trajectories (termed rollouts) under the current policy, starting from states in the offline dataset $\mathcal{D}_{\textrm{offline}}$.
Algorithm 1: Pseudocode for the base procedure used in offline model-based methods (MBPO, Janner et al. (2019)).
In summary:
[…]
Existing methods consider issues to be due to model errors and hence introduce dynamics uncertainty penalties. We instead reveal that the true source of issues is the edge-of-reach problem, and propose a value pessimism approach to directly address this.
Algorithm 1 Base model-based algorithm (MBPO)
+ Additions in existing methods and RAVL (ours)
1:  Require:Offline dataset $\mathcal{D}_{\textrm{offline}}$earned dynamics model $\widehat{M}$ = ($\widehat{T}$, $\widehat{R}$) trained on $\mathcal{D}_{\textrm{offline}}$ynamics uncertainty penalty added to $\widehat{M}$
2:  Specify:Rollout length $k\geq 1$, real data ratio $r\in[0,1]$
3:  Initialize:Replay buffer $\mathcal{D}_{\textrm{rollouts}}=\emptyset$, policy $\pi_{\theta}$, value function $Q_{\phi}$
4:  for $\textrm{epochs}=1,\dots$ do
5:     (Collect data)Starting from states in $\mathcal{D}_{\textrm{offline}}$, collect $k$-step rollouts in $\widehat{M}$ with $\pi_{\theta}$. Store data in $\mathcal{D}_{\textrm{rollouts}}$
6:     (Train agent)Train $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\textrm{rollouts}}\cup\mathcal{D}_{\textrm{offline}}$ mixed with ratio $r$ with value pessimism applied to the $Q$-function
7:  end for*Vanishes with an Oracle model (Table 1) as $\bm{\widehat{M}}$ replaced by true $\bm{M}$
3 Related Work
Overestimation bias in $Q$-learning. $Q$-learning is a popular basis for many off-policy RL algorithms, however, it has been shown to suffer from an overestimation bias even in the online case (van Hasselt et al., 2016). In online RL, Fujimoto et al. (2018) demonstrate taking the minimum over two independently trained $Q$-functions to be an effective solution, and extending this to larger $Q$-ensembles has been shown to be beneficial for further control of overestimation (Chen et al., 2021). An et al. (2021) similarly apply taking a minimum over a $Q$-ensemble in model-free offline RL to mitigate the more extreme pathological value overestimation caused by the out-of-sample problem.
Offline Model-Free Methods Model-free offline algorithms can broadly be divided into action constraint methods and value pessimism-based methods, which each represent a different approach to the out-of-sample problem. Action constraint methods (Kostrikov et al., 2022; Kumar et al., 2019; Fujimoto & Gu, 2021) aim to avoid using out-of-sample actions in the Bellman update by ensuring selected actions are close to the dataset behavior policy $\pi^{\beta}$. By contrast, value pessimism-based methods aim to directly regularize the value function such that it produces low-value estimates for out-of-sample state-actions (Kumar et al., 2020; Kostrikov et al., 2021; An et al., 2021), hence mitigating the overestimation mechanism described in Section 2. Of these, the most similar to our RAVL algorithm is EDAC (An et al., 2021) which applies value pessimism at any out-of-sample state-actions using minimization over a $Q$-ensemble. [^4] As discussed in Section 6, our RAVL approach can be viewed as reducing the required value pessimism compared to EDAC.
Offline Model-Based Methods As alluded to in Section 1, the prevailing (incorrect) view in the offline model-based literature is that the ability to collect additional data in the model effectively resolves the out-of-sample issue, as the agent is now able to try out any previously out-of-sample behavior. The primary concern stated for model-based methods is therefore to account for errors in the learned dynamics model to avoid model exploitation. A broad class of methods explicitly penalizes the reward by the estimated level of model uncertainty (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023; Lu et al., 2022; Ball et al., 2021), typically using variance over a dynamics ensemble. Rigter et al. (2022) aim to avoid model exploitation by setting up an adversarial two-player game between the policy and model. Critically, all previously described methods draw their motivation from dealing with dynamics model errors. Finally, most related to our method, COMBO (Yu et al., 2021), penalizes value estimates for state-actions outside model rollouts. However, similarly to Yu et al. (2020), COMBO is theoretically motivated by the assumption of infinite horizon model rollouts, which we later show overlooks serious implications.
4 The Edge-of-Reach Problem
In the following section, we describe the central insight of our paper: the edge-of-reach problem. We begin by presenting the surprising result which exposes an error in current understanding. Next, we present our hypothesis, including intuition, empirical evidence on the main D4RL benchmark (Fu et al., 2020), and theoretical proof of its effect on offline model-based training.
4.1 Surprising Result: Failure with the True Dynamics
Table 1: Surprising result: Oracle denotes replacing the learned dynamics in any dynamics-penalized method (MOPO, MOReL, MOBILE) with the true dynamics (since these collapse to the same procedure with no model error).5 This change surprisingly results in failure on the D4RL MuJoCo v2 datasets. For comparison with using Learned dynamics (with dynamics uncertainty penalization) we list MOPO’s results (as it is the weakest of the dynamics-penalized methods). Statistical significance highlighted (6 seeds).
Environment Learned Oracle
Halfcheetah random 38.5 35.2$\pm$3.4 $\downarrow$9%
medium 73.0 72.2$\pm$4.1 -0%
mixed 72.1 72.2$\pm$4.1 -0%
medexp 90.8 84.9$\pm$17.6 $\downarrow$7%
Hopper random 31.7 9.5$\pm$1.1 $\downarrow$70%
medium 62.8 $1$downarrow73%
mixed 103.5 71.5$\pm$32.2 $\downarrow$31%
medexp 81.6 6.1$\pm$6.8 $\downarrow$93%
Walker2d random 7.4 -0.4$\pm$0.2 $\downarrow$105%
medium 84.1 7.2$\pm$1.6 $\downarrow$91%
mixed 85.6 7.9$\pm$1.9 $\downarrow$91%
medexp 112.9 7.7$\pm$2.1 $\downarrow$93%
As described in Section 1, the current understanding in offline model-based RL is that, by adding a model to enable the collection of new data throughout training, the offline problem is effectively transformed into online learning. With this view, the key remaining difference compared to online RL is that: in online RL, data is collected with the true dynamics, while in offline model-based RL, data is collected in an approximate dynamics model. Existing model-based algorithms use this understanding, and as a result base their approaches on trying to mitigate the effect of dynamics model errors, for example by introducing reward or termination penalties based on estimates of dynamics uncertainty (Yu et al., 2020; Sun et al., 2023; Kidambi et al., 2020).
[…]
Figure 2: An illustration of how states that are sampled only in the final step of rollouts (i.e. as $s_{k}$) will appear only as nextstates in the model-based training distribution. These states will be used as targets in the Bellman update (see Equation 1), but will not appear as $s$ in the dataset, and hence never themselves be updated. This mismatch leads to a breakdown of Bellman-based learning.
4.2 The Edge-Of-Reach Hypothesis
Our observation in Section 4.1 reveals that issues in offline model-based RL cannot be explained by dynamics model errors - a widespread misconception. We therefore require a different understanding. We begin with the main intuition: in offline model-based RL, synthetic data $\mathcal{D}_{\textrm{rollouts}}$ is generated as short $k$-step rollouts starting from states in the original dataset $\mathcal{D}_{\textrm{offline}}$ (see Algorithm 1). Crucially, under this procedure, there can exist some edge-of-reach states which can be reached in the final rollout step, but which cannot - under any policy - be reached earlier (see Figure 1 red). Sampling these states triggers a breakdown in learning, since even with the ability to collect unlimited additional data, the agent is never able to reach them ‘in time’ to try actions from them and hence never receives feedback.
More concretely: Figure 2 illustrates how, due to being sampled only at the final step in a rollout, edge-of-reach states will be used for targets in Bellman-based learning but are never themselves updated. Thus, the out-of-sample action problem in model-free methods persists: since values at edge-of-reach states can be arbitrarily misestimated, incorrect values are then propagated over the entire state-action space. Moreover, when used with function approximation, the $\max$ operation in policy improvement can then exploit this misestimation, picking out the most heavily overestimated values (Levine et al., 2020). This results in extreme value overestimation, often with $Q$-values increasing unboundedly over training (as we see in Figure 3).
Edge-of-Reach Hypothesis: The procedure of limited horizon rollouts from a fixed dataset leads to the existence of edge-of-reach states at which the agent is unable to collect corrective feedback. This lack of corrective feedback triggers pathological value overestimation.
Do we expect this to actually happen in practice? Failure via this mode requires (a) existence of edge-of-reach states, and (b) such edge-of-reach states to be sampled. The typical combination of $k<<H$ along with a limited pool of starting states ($s\in\mathcal{D}_{\textrm{offline}}$) means the rollout distribution is unlikely to sufficiently cover the full state space $\mathcal{S}$, making (a) likely. Moreover, we observe pathological ‘edge-of-reach seeking’ behavior in which the agent appears to ‘seek out’ edge-of-reach states due to this being the source of overestimation (see Section 5), thus making (b) likely. In Appendix A, we give a more thorough and unified view of model-free and model-based offline RL, dividing the problem into independent conditions for states and actions.
4.3 Definitions and Formalization
We now formalize the intuitions presented in Section 4.2, and present theoretical grounding for the edge-of-reach problem and how errors can propagate to all transitions.
Definition 1 (Edge-of-reach states). Consider a deterministic transition model $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, rollout length $k$, and some distribution over starting states $\nu_{0}$. For some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$, rollouts are then generated according to $s_{0}\sim\nu_{0}(\cdot)$, $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim T(\cdot|s_{t},a_{t})$ for $t=0\dots k-1$, giving $(s_{0},a_{0},s_{1},\dots,s_{k})$. Let us use $\rho_{t,\pi}(s)$ to denote the marginal distributions over $s_{t}$.
We define a state $s\in S$ edge-of-reach with respect to ($T$, $k$, $\nu_{0}$) if: for $t=k$, $\exists\>\pi$ s.t. $\rho_{t,\pi}(s)>0$, but, for $t=1,\dots,k-1$ and $\forall\>\pi$, $\rho_{t,\pi}(s)=0$.
In our case, $\nu_{0}$ is the distribution of states in $\mathcal{D}_{\textrm{offline}}$. With stochastic transition models (e.g. Gaussian models), we may have the case that no state will truly have zero density, and we relax this definition slightly to $\rho_{t,\pi}(s)<\epsilon$ for some small $\epsilon$. The next proposition quantifies how errors in edge-of-reach states are propagated to preceding states.
Proposition 1 (Error propagation from edge-of-reach states). Consider a rollout of length $k$, $(s_{0},a_{0},s_{1},\dots,s_{k})$. Suppose that the state $s_{k}$ is edge-of-reach and the approximate value function $Q^{j}(s_{k},\pi(s_{k}))$ has error $\epsilon$. Then, standard value iteration will compound error $\gamma^{k-t}\epsilon$ to the estimates of $Q^{j+1}(s_{t},a_{t})$ for $t=1,\dots,k-1$. (Proof in Appendix D.)
In practice, model rollouts are sampled in minibatches and this effect will occur on average throughout training. This is an analogous statement to that of Kumar et al. (2019) in the offline model-free setting; however, its significance in the context of model-based methods has previously been overlooked. In the next section, we provide further evidence of this edge-of-reach problem being the predominant issue in offline model-based RL.
Table 2: Per-step rewards with the base offline model-based procedure (MBPO, see Algorithm 1). Rewards in model rollouts are close to those with the true dynamics, showing that model exploitation could not explain the subsequent value overestimation. Further, it indicates (as in Lu et al. (2022); Janner et al. (2019)) that the model is largely accurate for short rollouts and hence unlikely to be vulnerable to model exploitation.
Environment Model reward True reward
Hc-med 4.89$\pm$1.27 5.01$\pm$1.09
Hop-mixed 2.37$\pm$1.05 2.44$\pm$1.02
W2d-medexp 3.96$\pm$1.41 3.96$\pm$1.41
Figure 3: The base MBPO procedure results in poor performance (left) with an exponential increase in $Q$-values (right) on the D4RL benchmark. Mean and standard deviation given over 6 seeds. Approx $Q^{*}$ roughly indicates the $Q$-value for a normalized score of 100 (with $\gamma=0.99$). Results are shown for Walker2d-medexp, but we note similar trends across other D4RL datasets.
Figure 4: Experiments on the simple environment, illustrating the edge-of-reach problem and potential solutions. (a) Reward function, (b) final (failed) policy with naïve application of SAC, (c) final (successful) policy with patching in oracle $Q$-values for edge-of-reach states, (d) final (successful) policy with RAVL, (e) returns evaluated over training, (f) mean $Q$-values evaluated over training.
4.4 Evidence for the Edge-of-Reach Problem
There are two potential issues in offline model-based RL: (1) dynamics model errors and subsequent model exploitation, and (2) the edge-of-reach problem and subsequent pathological value overestimation. We ask: Which is the true source of the issues observed in practice? While (1) is stated as the sole issue and hence the motivation in prior methods, the results in Table 1 are strongly at odds with this explanation. Furthermore, in Table 2 we check and find that the difference between the rewards predicted by the model vs the true rewards is insignificant, which means that (1) fails to explain the value explosion in Figure 3. Instead, edge-of-reach-induced pathological overestimation mechanism of (2) exactly predicts this value explosion, with the observations very closely resembling those of the analogous model-free out-of-sample problem (Kumar et al., 2019). Furthermore, (2) is consistent with the ‘surprising result’ observations in Table 1.
5 Analysis with a Simple Environment
In the previous section, we provided a plausible theoretical explanation and intuition for unexpected failure observed in Section 4.1. In this section, we construct a simple environment and use careful analysis to empirically confirm this hypothesis. We first show that naïve application of actor-critic algorithms results in exploding $Q$-values and failure to learn despite using the true dynamics model. Next, we verify that edge-of-reach states are the source of this problem by showing that correcting value estimates only at these states is sufficient to resolve the problem.
5.1 Setup
[…]
The only difference here compared to online RL is that we use a data collection horizon of $k=10$ such that $k<H$. This small change results in the existence of edge-of-reach states (those between the red and orange boxes in Figure 4), and hence captures the problematic component of the MBPO procedure (see Algorithm 1). The procedure is therefore training SAC in the true environment with limited length trajectories, exactly as is done in Experiment 4.1. Note: There is no notion of an approximate dynamics model, as we are demonstrating that the issues are not caused by dynamics model errors, as is currently assumed. We instead demonstrate that it is the existence of edge-of-reach states (due to the use of $k<H$ as per the standard MBPO procedure) that triggers pathological value overestimation.
5.2 Observing Pathological Value Overestimation
Since the only difference between this setup and online RL is the use of trajectory collection horizon $k<H$ (as in MBPO), we may expect SAC to learn a sensible policy. However, we observe that SAC fails, with $Q$-values growing unboundedly over training (see blue Figure 4(c,d)). Looking at the rollout trajectories sampled over training (see Figure 7) we see the following: (Before 25 epochs) Initially, the agent follows the reward function, and performance increases. (Between 25 and 160 epochs) Value misestimation takes over, and the policy begins to aim toward unobserved state-actions (since their values can be misestimated and hence overestimated), and performance subsequently decreases. (After 160 epochs) This ‘edge-of-reach seeking’ behavior compounds with each epoch, leading to the agent eventually reaching edge-of-reach states. From this point onwards, the agent samples edge-of-reach states at which it never receives any corrective feedback. The consequent pathological value overestimation results in a complete collapse in performance. In Figure 4(b) we visualize the final policy and see that it completely ignores the reward function. Instead, it aims towards an arbitrary edge-of-reach state with the highest (heavily overestimated) $Q$-value.
5.3 Verifying the Hypothesis Using Value Patching
Our hypothesis is that the source of this problem is value misestimation at edge-of-reach states. We verify this with our SAC-OraclePatch experiments, in which we replace a subset of the value estimates in the Bellman targets with their true values. In yellow in Figure 4 we demonstrate that applying patching solely for edge-of-reach states is sufficient to completely solve the problem. This is particularly compelling as in practice only a very small proportion of states had correct values patched in (0.4% of states over training). Of course, patching in true values is not possible in practice, thus in the next section we develop a practical approximation to this. We show in Section 7.1 that our proposed method, RAVL (see green in Figure 4), has a very similar effect to that of the ideal SAC-OraclePatch intervention and completely resolves the edge-of-reach problem.
Table 3: A comprehensive evaluation of RAVL over the standard D4RL MuJoCo benchmark. We show the mean and standard deviation of the final performance averaged over 6 seeds. Our simple approach matches the state-of-the-art without any explicit dynamics penalization.
Model-Free Model-Based
Environment BC CQL EDAC MOPO COMBO RAMBO MOBILE RAVL (Ours)
Halfcheetah medium 43.2 46.9 65.9 73.0 54.2 77.9 74.6 78.7$\pm$2.0
mixed 37.6 45.3 61.3 72.1 55.1 68.7 71.7 74.9$\pm$2.0
medexp 44.0 95.0 106.3 90.8 90.0 95.4 108.2 102.1$\pm$8.0
Hopper medium 54.1 61.9 101.6 62.8 97.2 87.0 106.6 90.6$\pm$11.9
mixed 16.6 86.3 101.0 103.5 89.5 99.5 103.9 103.1$\pm$1.1
medexp 53.9 96.9 110.7 81.6 111.1 88.2 112.6 110.1$\pm$1.8
Walker2d medium 70.9 79.5 92.5 84.1 81.9 84.9 87.7 86.3$\pm$1.6
mixed 20.3 76.8 87.1 85.6 56.0 89.2 89.9 83.0$\pm$2.5
medexp 90.1 109.1 114.7 112.9 103.3 56.7 115.2 115.5$\pm$2.4
6 RAVL: Reach-Aware Value Learning
As verified in Section 5.3, the edge-of-reach problem stems from value overestimation at edge-of-reach states. To resolve this issue, we therefore need to (A) detect and (B) prevent value overestimation at edge-of-reach states.
(A) Detecting edge-of-reach states: As illustrated in Figure 2, edge-of-reach states are those that remain outside the distribution of states being updated, i.e. the outside training distribution of the $Q$-function, $\mathcal{D}_{\textrm{rollouts}}$. A natural solution to detecting out-of-distribution (OOD) inputs is to use a deep ensemble (Lakshminarayanan et al., 2017). As such, high variance over an ensemble of $Q$-functions can be used to detect edge-of-reach states. We demonstrate that this is effective in Figure 5. Note this is different to related model-free methods since the distribution being considered is that of $\mathcal{D}_{\textrm{rollouts}}$ rather than $\mathcal{D}_{\textrm{offline}}$.
(B) Preventing overestimation at these states: Once we have detected edge-of-reach states, we may simply apply value pessimism methods from the offline model-free literature to these states. Our choice of an ensemble for part (A) conveniently allows us to simply minimize over an ensemble of $Q$-functions, which effectively penalizes values based on ensemble variance.
Our resulting proposal is Reach-Aware Value Learning (RAVL). Concretely, RAVL simply involves training an ensemble of $N_{\textrm{critic}}$ $Q$-functions similarly to EDAC (An et al., 2021), indexed by $n$, and modifying Equation 1 to:
$\small\begin{split}Q_{n}^{j+1}\leftarrow\underset{Q_{n}}{\arg\min}\;\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D},a^{\prime}\sim\pi^{j}(\cdot\mid s^{\prime})}[\qquad\qquad\>\\ (\underbrace{Q_{n}(s,a)}_{\text{\tiny input}}-\underbrace{[r+\gamma\min_{n=1,\dots,N_{\textrm{critic}}}Q_{n}^{j}(s^{\prime},a^{\prime})]}_{\text{\tiny Bellmantarget}})^{2}\>]\end{split}$ (2)
As is standard in offline model-based RL, we optimize this objective using MBPO (Janner et al., 2019) and SAC (Haarnoja et al., 2018) as in Algorithm 1. RAVL is therefore a solution to offline model-based RL, which relies solely on value pessimism ideas from the model-free literature. This is in stark contrast to typical methods which use dynamics uncertainty penalization and therefore also provides further evidence that the edge-of-reach problem is the true source of issues.
RAVL represents combining model-based training with the current SOTA model-free algorithm EDAC. The addition of the model compared to EDAC can be seen as relaxing the value pessimism requirements from any state-actions OOD with respect to $\mathcal{D}_{\textrm{offline}}$, to those OOD with respect to $\mathcal{D}_{\textrm{rollouts}}$. This perspective highlights our algorithm’s simplicity, with the approach being made possible due to our corrected understanding of the true issues involved in model-based offline RL. Our subsequently unified view of model-based and model-free approaches has allowed us to transfer ideas from the model-free literature for dealing with the out-of-sample problem and directly apply them to solve the edge-of-reach issue in the model-based setting.
7 Empirical Evaluation
In this section, we first evaluate RAVL on the simple environment from Section 5. Next, we show that RAVL scales, achieving strong performance on the standard D4RL benchmark and a new state-of-the-art on challenging pixel-based V-D4RL tasks. Finally, we compare RAVL to prior dynamics uncertainty penalized methods and elucidate why they may work despite not explicitly addressing the edge-of-reach problem. We provide full hyperparameters and a discussion of run-time in Section C.2.
7.1 Evaluation on the Simple Environment
We first test RAVL on the simple environment from Section 5 (see green in Figure 4). We observe that it achieves the same result as the theoretically optimal but practically impossible method of patching in the true values at edge-of-reach states, SAC-OraclePatch. The $Q$-values are stabilized and match those of SAC-OraclePatch. Moreover, in Figure 5, we demonstrate that RAVL can precisely detect edge-of-reach states as they have significantly higher ensemble variance than within-reach states, as expected.
Figure 5: RAVL’s effective penalty of variance over the $Q$-value ensemble on the environment in Section 5. As expected, values at edge-of-reach states have significantly higher variance than for within-reach states, meaning, exactly as desired, RAVL is able to detect and apply a value pessimism penalty at edge-of-reach states.
7.2 Evaluation on the D4RL Benchmark
Next, we show that RAVL scales to the standard offline RL benchmark, D4RL (Fu et al., 2020). We consider the MuJoCo (Todorov et al., 2012) v2 datasets. Figure 3 shows RAVL successfully stabilizes training, and Table 3 shows that we can match the performance of the current top model-based method, MOBILE (Sun et al., 2023), without the need for any explicit dynamics penalization. This is beneficial in terms of runtime, as MOBILE requires multiple transition samples to calculate its penalty. Compared to EDAC (An et al., 2021), the model-free equivalent of our algorithm, RAVL achieves much higher and state-of-the-art performance on the Halfcheetah mixed and medium datasets. This demonstrates the clear benefit of generating additional model-based samples. Our approach is far stronger than COMBO (Yu et al., 2021) which introduced a less effective form of value conservatism into model-based methods.
7.3 Evaluation on Pixel-Based tasks
The insights from RAVL also lead to improvements in performance on the challenging pixel-based benchmark V-D4RL (Lu et al., 2023). This uses the Offline DreamerV2 algorithm which trains on trajectories in a learned latent space, thus demonstrating that the edge-of-reach problem is more general and widespread. We observe in Table 4 that RAVL gives a strong boost in performance (+20.3% overall) on the medium and medexp level datasets, while not helping in the more diverse random and mixed datasets. This observation fits with our intuition of the edge-of-reach problem: the medium and medexp level datasets are likely to have less coverage of the state space, and thus we would expect them to suffer more from edge-of-reach issues and the ‘edge-of-reach seeking behavior’ demonstrated in Section 5.2. We note also that these results are without the ensemble diversity regularizer from EDAC (An et al., 2021) (used on the D4RL benchmark), which we anticipate may further increase performance.
Table 4: We show that RAVL extends to the challenging pixel-based V-D4RL benchmark. The change in mean performance is $+20.3\%$. Mean and standard deviation given over 6 seeds.
Environment Offline DV2 RAVL (Ours)
Walker-walk medium 34.1$\pm$19.7 48.6$\pm$8.7
medexp 43.9$\pm$34.4 47.6$\pm$18.3
Cheetah-run medium 17.2$\pm$3.5 18.3$\pm$4.3
medexp 10.4$\pm$3.5 12.6$\pm$4.4
7.4 Reinterpretation of Prior Methods
Prior model-based methods were unaware of the edge-of-reach problem, instead being motivated solely by mitigating issues due to model errors. However, our investigation strongly suggests that the edge-of-reach problem is likely to be the true main underlying issue on the D4RL benchmark (see Section 4.4). This leads us to ask: What allows prior model error-motivated methods to work despite ignoring the critical edge-of-reach problem? We find that, while ostensibly to address model errors, existing dynamics penalties may accidentally help to address the edge-of-reach problem, as we observe a positive correlation between the penalties used in dynamics uncertainty methods and RAVL’s effective penalty of value ensemble variance (see Figure 6 in Section B.1). This may be expected, as dynamics uncertainty will naturally be higher further away from $\mathcal{D}_{\textrm{offline}}$, which is also where edge-of-reach states are more likely to lie. We additionally note that, since the edge-of-reach problem is orthogonal to dynamics misestimation, RAVL could be combined with appropriate model uncertainty penalization (Yu et al., 2020; Sun et al., 2023; Lu et al., 2022) for environments with less accurate models.
8 Conclusion
In this paper, we expose a major misunderstanding in current model-based offline RL. We provide strong evidence that - contrary to common understanding - the true source of issues experienced by current methods is the edge-of-reach problem, a previously unnoticed issue whereby the standard offline model-based RL procedure leads to the existence of a set of edge-of-reach states that triggers pathological value misestimation and collapse of learning. Exposing this reveals a strong connection between model-free and model-based offline RL, which allows us to unify our understanding of these two areas. Based on this new unified perspective, we are able to transfer ideas from model-free literature and directly apply them to solve the edge-of-reach problem in model-based methods. Our resulting practical algorithm RAVL, is simple, robust, and achieves strong performance across both proprioceptive and pixel-based benchmarks. We open-source our code and hope that the insights presented here may help further ideas to be transferred between model-free and model-based methods. Finally, a more complete theoretical picture could be obtained by transferring finite horizon rollout theory from online model-based reinforcement learning to the offline setting (Janner et al., 2019).
Statement of Broader Impact
This paper presents work whose goal is to advance the field of offline reinforcement learning. Since we evaluate on simulated control problems in this paper, we do not expect any specific ethical concerns to arise. However, when offline reinforcement learning is deployed in the real world, it is important to follow responsible ethical principles in ensuring the trustworthiness and reliability of our systems.
Acknowledgements
Anya Sims and Cong Lu are funded by the Engineering and Physical Sciences Research Council (EPSRC). The authors would like to thank the anonymous Agent Learning in Open-Endedness Workshop at NeurIPS 2023 and ICLR 2024 reviewers for positive and constructive feedback which helped to improve the paper. We would also like to thank Philip J. Ball and Shimon Whiteson for helpful feedback and discussions on earlier drafts of this work.
Supplementary Material
Appendix A A Unified Perspective of Model-Based and Model-Free RL
We supplement the discussion in Section 4.2 with a more thorough comparison of the out-of-sample and edge-of-reach problems, including how they relate to model-free and model-based approaches.
A.1 Definitions
Consider a dataset of transition tuples $\mathcal{D}=\{(s_{i},a_{i},r_{i},s^{\prime}_{i},d_{i})\}_{i=1,\dots,N}$ collected according to some dataset policy $\pi^{\mathcal{D}}(\cdot|s)$. Compared to Section 2, we include the addition of a done indicator $d_{i}$, where $d_{i}=1$ indicates episode termination (and $d_{i}=0$ otherwise). Transition tuples thus consist of state, action, reward, nextstate, done. Consider the marginal distribution over state-actions $\rho^{\mathcal{D}}_{s,a}(\cdot,\cdot)$, over states $\rho^{\mathcal{D}}_{s}(\cdot)$, and conditional action distribution $\rho^{\mathcal{D}}_{a|s}(\cdot|s)$. Note that $\rho^{\mathcal{D}}_{a|s}(\cdot|s)=\pi^{\mathcal{D}}(\cdot|s)$. We abbreviate $x$ is in distribution with respect to $\rho$ as $x\in^{\textrm{dist}}\rho$.
A.2 $Q$-Learning Conditions
As described in Section 2, given some policy $\pi$, we can attempt to learn the corresponding $Q$-function with the following iterative process:
[…]
$Q$-learning relies on bootstrapping, hence to be successful we need to be able to learn accurate estimates of the Bellman targets for all $(s,a)$ inputs. Bootstrapped estimates of $Q(s^{\prime},a^{\prime})$ are used in the targets whenever $d\neq 1$. Therefore, for all $(s^{\prime},a^{\prime})$, we require:
  • Combined state-action condition: $(s^{\prime},a^{\prime})\in^{\textrm{dist}}\rho^{\mathcal{D}}_{s,a}$ or $d=1$.
In the main paper, we use this combined state-action perspective for simplicity, however, we can equivalently divide this state-action condition into independent requirements on the state and action as follows:
  • State condition: $s^{\prime}\in^{\textrm{dist}}\rho^{\mathcal{D}}_{s}$ or $d=1$,
  • Action condition: $a^{\prime}\in^{\textrm{dist}}\rho^{\mathcal{D}}_{a|s}(s^{\prime})$ (given the above condition is met and $d\neq 1$).
Informally, the state condition may be violated if $\mathcal{D}$ consists of partial or truncated trajectories, and the action condition may be violated if there is a significant distribution shift between $\pi^{\mathcal{D}}$ and $\pi$.
A.3 Comparison Between Model-Free and Model-Based Methods
Table 5: A summary of the comparison between model-free and model-based offline RL in relation to the conditions on $Q$-learning.
Action condition violation State condition violation Main source of issues
Model-free Possible, common in practice Possible, but uncommon in practice $\rightarrow$ Action condition violation
Model-based Not possible Possible, common in practice $\rightarrow$ State condition violation
In offline model-free RL, $\mathcal{D}=\mathcal{D}_{\textrm{offline}}$, with $\pi^{\mathcal{D}}=\pi^{\beta}$. For the settings we consider, $\mathcal{D}_{\textrm{offline}}$ consists of full trajectories and therefore will not violate the state condition. However, this may happen in a more general setting with $\mathcal{D}_{\textrm{offline}}$ containing truncated trajectories. By contrast, the mismatch between $\pi$ (used to sample $a^{\prime}$ in $Q$-learning) and $\pi^{\beta}$ (used to sample $a$ in the dataset $\mathcal{D}_{\textrm{offline}}$) often does lead to significant violation of the action condition. This exacerbates the overestimation bias in $Q$-learning (see Section 3), and can result in pathological training dynamics and $Q$-value explosion over training (Kumar et al., 2019).
On the other hand, in offline model-based RL, the dataset $\mathcal{D}=\mathcal{D}_{\textrm{rollouts}}$ is collected on-policy according to the current (or recent) policy such that $\pi^{\mathcal{D}}\approx\pi$. This minimal mismatch between $\pi^{\mathcal{D}}$ and $\pi$ means the action condition is not violated and can be considered to be resolved due to the collection of additional data. However, the procedure of generating the data $\mathcal{D}=\mathcal{D}_{\textrm{rollouts}}$ can be seen to significantly exacerbate the state condition problem, as the use of short truncated-horizon trajectories means the resulting dataset $\mathcal{D}_{\textrm{rollouts}}$ is likely to violate the state condition. Due to lack of exploration, certain states may temporarily violate the state condition. Our paper then considers the pathological case of edge-of-reach states, which will always violate the state condition.
Appendix B Additional Figures
B.1 Comparison to Prior Approaches
In Figure 6, we plot the dynamics uncertainty-based penalty used in MOPO (Yu et al., 2020) against the effective penalty in RAVL of variance of the value ensemble. The positive correlation between MOPO’s penalty and the edge-of-reach states targeting penalty of RAVL helps to explain why prior methods may work despite not considering the crucial edge-of-reach problem.
Figure 6: We find that the dynamics uncertainty-based penalty used in MOPO (Yu et al., 2020) is positively correlated with the variance of the value ensemble of RAVL, suggesting prior methods may unintentionally address the edge-of-reach problem. Pearson correlation coefficients are 0.49, 0.43, and 0.27 for Hopper-mixed, Walker2d-medexp, and Halfcheetah-medium respectively.
B.2 Visualizations of Rollouts Through Training
In Figure 7, we provide a visualization of the rollouts sampled over training in the simple environment for each of the algorithms analyzed in Figure 4 (see Section 5). This accompanies the discussion of the behavior of SAC during training in Section 5.2.
Figure 7: A visualization of the rollouts sampled over training on the simple environment in Section 5. We note the pathological behavior of the baseline SAC, and the success of the ideal intervention SAC-OraclePatch, and our practically realizable method RAVL.
Appendix C Implementation Details
In this section, we give full implementation details for RAVL.
C.1 Algorithm
As is typical in offline model-based reinforcement learning (Yu et al., 2020; Lu et al., 2022; Kidambi et al., 2020; Sun et al., 2023), we use short MBPO (Janner et al., 2019) style rollouts. The agent is trained using the EDAC (An et al., 2021) losses, a conservative ensemble-based algorithm based on SAC (Haarnoja et al., 2018). Our transition model follows the standard setup in model-based offline algorithms, being realized as a deep ensemble (Chua et al., 2018) and trained via maximum likelihood estimation. This is summarized in the pseudocode for offline model-based methods given in Algorithm 1.
C.2 Hyperparameters
For the D4RL (Fu et al., 2020) MuJoCo datasets presented in Section 7.2, we sweep over the following hyperparameters and list the optimal choices in Table 6.
  • (EDAC) Number of $Q$-ensemble elements $N_{\textrm{critic}}$, in the range $\{10,50\}$
  • (EDAC) Ensemble diversity weight $\eta$, in the range $\{1,10,100\}$
  • (MBPO) Model rollout length $k$, in the range $\{1,5\}$
  • (MBPO) Real-to-synthetic data ratio $r$, in the range $\{0.05,0.5\}$
The remaining model-based and agent hyperparameters are given in Table 7. Almost all environments use a small $N_{\textrm{critic}}=10$ similar to An et al. (2021), with only the Hopper-medexp dataset needing $N_{\textrm{critic}}=30$. Our implementation is based on the Clean Offline Reinforcement Learning (CORL, Tarasov et al. (2022)) repository, released at https://github.com/tinkoff-ai/CORL under an Apache-2.0 license. Our algorithm takes approximately 1 day to run on the standard D4RL datasets using an A100 GPU for the full number of epochs.
Table 6: Variable hyperparameters for RAVL used in D4RL MuJoCo locomotion tasks.
Environment $\mathbf{N_{critic}}$ $\bm{\eta}$ $\mathbf{k}$ $\mathbf{r}$
HalfCheetah medium 10 1 5 0.05
mixed 10 100 5 0.05
medexp 10 1 5 0.5
Hopper medium 10 100 1 0.5
mixed 10 10 1 0.5
medexp 50 100 1 0.5
Walker2d medium 10 10 1 0.5
mixed 10 10 1 0.5
medexp 10 1 1 0.5
Table 7: Fixed hyperparameters for RAVL used in D4RL MuJoCo locomotion tasks.
Parameter Value
epochs 3,000 for medexp; 1,000 for rest
gamma 0.99
learning rate $3\times 10^{-4}$
batch size 256
buffer retain epochs 5
number of rollouts 50,000
C.3 Pixel-Based Hyperparameters
For the V-D4RL (Lu et al., 2023) DeepMind Control Suite datasets presented in Section 7.3, we use the default hyperparameters for the Offline DV2 algorithm, which are given in Table 8. We found keeping the uncertainty weight $\lambda=10$ improved performance over $\lambda=0$ which shows RAVL can be combined with dynamics penalty-based methods.
Table 8: Hyperparameters for RAVL used in V-D4RL DeepMind Control Suite tasks.
Parameter Value
ensemble member count (K) 7
imagination horizon (H) 5
batch size 64
sequence length (L) 50
action repeat 2
observation size [64, 64]
discount ($\gamma$) 0.99
optimizer Adam
learning rate {model = $3\times 10^{-4}$, actor-critic = $8\times 10^{-5}$}
model training epochs 800
agent training epochs 2,400
uncertainty penalty mean disagreement
uncertainty weight ($\lambda$) 10
number of critics ($\textrm{N}_{\textrm{critic}}$) 5
We swept over $\{2,5,20\}$ for $\textrm{N}_{\textrm{critic}}$ but found a single value of $5$ worked well for all environments we consider.
Appendix D Proof of Error Propagation Result
In this section, we provide a proof of Proposition 1. Our proof follows analogous logic to the error propagation result of Kumar et al. (2019).
Proof.
Let us denote $Q^{\ast}$ as the optimal value function, $\zeta_{j}(s,a)=|Q_{j}(s,a)-Q^{\ast}(s,a)|$ the error at iteration $j$ of Q-Learning, and $\delta_{j}(s,a)=|Q_{j}(s,a)-\mathcal{T}Q_{j-1}(s,a)|$ the current Bellman error. Then first considering the $t=k-1$ case,
$\zeta_{j}(s_{t},a_{t})$ $=|Q_{j}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (4)
$=|Q_{j}(s_{t},a_{t})-\mathcal{T}Q_{j-1}(s_{t},a_{t})+\mathcal{T}Q_{j-1}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (5)
$\leq|Q_{j}(s_{t},a_{t})-\mathcal{T}Q_{j-1}(s_{t},a_{t})|+|\mathcal{T}Q_{j-1}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (6)
$=\delta_{j}(s_{t},a_{t})+\gamma\zeta_{j-1}(s_{t+1},a_{t+1})$ (7)
$=\delta_{j}(s_{t},a_{t})+\gamma\epsilon$ (8)
Thus the errors at edge-of-reach states are discounted and then compounded with new errors at $Q^{j}(s_{k-1},a_{k-1})$. For $t<k-1$, the result follows from repeated application of Equation 7 along the rollout.
Appendix E Summary of Setups Used in Comparisons
Throughout the paper, we compare several different setups in order to identify the true underlying issues in model-based offline RL. We provide a summary of them in Table 9. More comprehensive descriptions of each are given in the main text and in relevant table and figure captions.
Table 9: We summarize the various setups used for comparisons throughout the paper. ‘*’ denotes application to the simple environment (see Section 5). All methods use $k$-step rollouts from the offline dataset (or from a fixed starting state distribution in the case of the simple environment). Use of the MOPO penalty refers to implementing the method in Yu et al. (2020).
Dynamics Model Agent Figures and Tables
Type Penalty $N_{\textrm{critic}}$ $\eta$
Learned Ensemble ✓(MOPO) $2$ 00 Table 1, Figure 6
Oracle True n/a $2$ 00 Table 1
MBPO (Offline) Ensemble $2$ 00 Tables 2, Figure 3
RAVL (Ours) Ensemble $>2$ $>0$ Tables 2, 3, Figure 3
SAC (Truncated Horizon)* True n/a $2$ 00 Figures 4, 7
RAVL (Ours)* True n/a $>2$ $>0$ Figures 4, 5, 7

### v2

@@
 Abstract
-Offline reinforcement learning aims to enable agents to be trained from pre-collected datasets, however, this comes with the added challenge of estimating the value of behavior not covered in the dataset.
-Model-based methods offer a solution by allowing agents to collect additional synthetic data via rollouts in a learned dynamics model.
-The prevailing theoretical understanding is that this can then be viewed as online reinforcement learning in an approximate dynamics model, and any remaining gap is therefore assumed to be due to the imperfect dynamics model.
-Surprisingly, however, we find that if the learned dynamics model is replaced by the true error-free dynamics, existing model-based methods completely fail.
-This reveals a major misconception.
-Our subsequent investigation finds that the general procedure used in model-based algorithms results in the existence of a set of edge-of-reach states which trigger pathological value overestimation and collapse in Bellman-based algorithms.
+Offline reinforcement learning (RL) aims to train agents from pre-collected datasets.
+However, this comes with the added challenge of estimating the value of behaviors not covered in the dataset.
+Model-based methods offer a potential solution by training an approximate dynamics model, which then allows collection of additional synthetic data via rollouts in this model.
+The prevailing theory treats this approach as online RL in an approximate dynamics model, and any remaining performance gap is therefore understood as being due to dynamics model errors.
+In this paper, we analyze this assumption and investigate how popular algorithms perform as the learned dynamics model is improved.
+In contrast to both intuition and theory, if the learned dynamics model is replaced by the true error-free dynamics, existing model-based methods completely fail.
+This reveals a key oversight: The theoretical foundations assume sampling of full horizon rollouts in the learned dynamics model; however, in practice, the number of model-rollout steps is aggressively reduced to prevent accumulating errors.
+We show that this truncation of rollouts results in a set of edge-of-reach states at which we are effectively “bootstrapping from the void.” This triggers pathological value overestimation and complete performance collapse.
 We term this the edge-of-reach problem.
-Based on this, we fill some gaps in existing theory and also explain how prior model-based methods are inadvertently addressing the true underlying edge-of-reach problem.
-Finally, we propose Reach-Aware Value Learning (RAVL), a simple and robust method that directly addresses the edge-of-reach problem and achieves strong performance across both proprioceptive and pixel-based benchmarks.
-Offline Reinforcement Learning, Model-Based Reinforcement Learning
+Based on this new insight, we fill important gaps in existing theory, and reveal how prior model-based methods are primarily addressing the edge-of-reach problem, rather than model-inaccuracy as claimed.
+Finally, we propose Reach-Aware Value Learning (RAVL), a simple and robust method that directly addresses the edge-of-reach problem and hence - unlike existing methods - does not fail as the dynamics model is improved.
+Since world models will inevitably improve, we believe this is a key step towards future-proofing offline RL.
+[^1] Our code is open-sourced at: github.com/anyasims/edge-of-reach.
 1 Introduction
-Figure 1: Boundaries of the states reachable at each timestep in the $k$-step rollouts from $\mathcal{D}_{\textrm{offline}}$ as used in offline model-based RL.
-In this paper we investigate a hitherto unrecognized pathology in offline model-based methods whereby edge-of-reach states - those only reachable at the final step of model rollouts (in red) - can trigger value overestimation and collapse of training.
-Standard online reinforcement learning (RL) requires the collection of large amounts of on-policy data throughout training.
-This can be both costly and unsafe, and hence represents a barrier against applying RL in domains such as healthcare (Tang & Wiens, 2021; Shiranthika et al., 2022) or robotics (Kumar et al., 2021; Chebotar et al., 2021; Ball et al., 2021), and also against scaling RL to more complex problems.
-Offline RL (Levine et al., 2020; Ernst et al., 2005) aims to remove this need for online data collection by enabling agents to be trained on pre-collected datasets.
-The use of large pre-existing datasets has driven significant recent advances in supervised learning (Brown et al., 2020), and one hope is that offline RL may provide a route for similar scaling in RL (Kumar et al., 2023).Code open-sourced at: github.com/anyasims/edge-of-reach.
-The central challenge in offline RL is dealing with value estimation for counterfactual actions not present in the dataset, termed the out-of-sample action problem (Kostrikov et al., 2022).
-[^1] Also referred to as the out-of-distribution action (Kumar et al., 2019) or action distribution shift problem (Kumar et al., 2020).
-A naïve approach results in inaccurate values at out-of-sample state-actions being propagated through the entire state-action space, leading to pathological training dynamics often exhibited as extreme value overestimation (Kumar et al., 2019).
-There have been many proposals to resolve this, with methods largely falling into one of two categories: model-free (Kostrikov et al., 2022; Kumar et al., 2020; Fujimoto et al., 2019; Kumar et al., 2019; An et al., 2021; Fujimoto & Gu, 2021) or model-based (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023; Lu et al., 2022).
-Model-free methods typically address this issue by applying a form of conservatism or constraint to avoid using out-of-sample actions in the Bellman update.
-In contrast, model-based methods, which are the focus of this paper, generally aim to resolve the issue by using the offline dataset to fit an approximate dynamics model (Sutton, 1991; Janner et al., 2019), and then allowing the agent to collect additional data in this learned dynamics model through fixed horizon rollouts from the original offline dataset (the general MBPO procedure (Janner et al., 2019), see Algorithm 1).
-The current understanding is that this converts the offline learning problem into online RL in an approximate dynamics model, thus solving the out-of-sample problem by allowing the agent to collect corrective feedback for any misestimated values.
-With this understanding, model-based methods typically attribute any remaining difficulties to errors in the learned dynamics model, and algorithms are subsequently motivated by avoiding exploitation of model errors using some notion of dynamics uncertainty.
-This understanding of model-based approaches naturally implies that reducing dynamics model errors should improve performance.
-However, we begin with a surprising demonstration that existing offline model-based methods completely fail when trained with the true error-free dynamics model (see Table 1 below).
-This exposes a significant misconception in the current understanding of model-based offline RL and prompts us to reevaluate our understanding of how prior methods work.
-On investigation, we find that the use of short rollouts in these algorithms (as opposed to full horizon trajectories as collected in online RL) causes a critical and previously unnoticed problem.
-More specifically, it leads to the existence of states which, under any policy, can only be reached in the final rollout step (see red in Figure 1).
-These edge-of-reach states appear in the training distribution as nextstates, and are hence used in Bellman updates, but even with the ability to collect unlimited additional data, the agent is never able to observe the outcome of actions from them.
-The result is pathological overestimation of values, similar to that seen in model-free methods.
-We term this the edge-of-reach problem.
-Thus, contrary to common understanding, the out-of-sample problem central in model-free methods is not resolved by a model-based approach.
-This edge-of-reach problem reveals strong connections between the two previously disjoint subfields of offline model-free and model-based RL.
-With this new understanding, we propose Reach-Aware Value Learning (RAVL), a robust scalable solution that takes an existing solution to the central out-of-sample problem in model-free literature, and directly applies it to the edge-of-reach problem in model-based methods (both proprioceptive and pixel-based).
-We first analyze RAVL on simple toy environments, and then show that it scales to the standard D4RL (Fu et al., 2020) and the V-D4RL (Lu et al., 2023) benchmarks.
-Furthermore, we anticipate that this unified understanding of model-based and model-free algorithms may pave the way for porting further strengths of existing model-free algorithms into model-based approaches.
-Code is open-sourced at: github.com/anyasims/edge-of-reach
-To summarize, our contributions are as follows:
-  • We expose a significant misconception in the current understanding of offline model-based RL (Section 4.1).
-This leads to our hypothesis and definition of the edge-of-reach problem (Section 4.2).
-We provide evidence that this is the true underlying source of issues on the standard D4RL (Fu et al., 2020) benchmark (Section 4.4), before then constructing a simple environment which allows us to demonstrate and verify this hypothesis (Section 5).
-  • Based on this, we propose Reach-Aware Value Learning (RAVL), a natural solution to the edge-of-reach problem (Section 6).
-We extensively evaluate RAVL across the D4RL and pixel-based V-D4RL (Lu et al., 2023) benchmarks (Section 7), including careful analysis into what RAVL is learning, demonstrating that it directly targets the edge-of-reach problem as intended (Section 7.1) and achieves state-of-the-art performance.
-  • Finally, we give a more accurate explanation of the efficacy of prior methods, showing they unintentionally address the edge-of-reach problem (Section 7.4).
-[^2] Throughout the paper, please refer to Table 9 for the naming conventions used in experiments.
+Standard online reinforcement learning (RL) requires collecting large amounts of on-policy data.
+This can be both costly and unsafe, and hence represents a significant barrier against applying RL in domains such as healthcare [32, 28] or robotics [20, 4, 2], and also against scaling RL to more complex problems.
+Offline RL [23, 6] aims to remove this need for online data collection by enabling agents to be trained on pre-collected datasets.
+One hope [21] is that it may facilitate advances in RL similar to those driven by the use of large pre-existing datasets in supervised learning [3].
+[…]
+The central challenge in offline RL is estimating the value of actions not present in the dataset, known as the out-of-sample action problem [17].
+[^2] This is also referred to as the out-of-distribution action [18] or action distribution shift problem [19].
+A naïve approach results in extreme value overestimation due to bootstrapping using inaccurate values at out-of-sample state-actions [18].
+There have been many proposals to resolve this, with methods largely falling into one of two categories: model-free [17, 19, 10, 18, 1, 8] or model-based [36, 14, 29, 24].
+Model-free methods typically address the out-of-sample action problem by applying a form of conservatism or constraint to avoid using out-of-sample actions in the Bellman update.
+In contrast, the solution proposed by model-based methods is to allow the collection of additional data at any previously out-of-sample actions.
+This is done by first training an approximate dynamics model on the offline dataset [30, 13], and then allowing the agent to collect additional synthetic data in this model via $k$-step rollouts (see Algorithm 1).
+The prevailing understanding is that this can be viewed as online RL in an approximate dynamics model, with the instruction being to then simply “run any RL algorithm on $\widehat{M}$ until convergence”[36], where $\widehat{M}$ is the learned dynamics model with some form of dynamics uncertainty penalty.
+Existing methods propose various forms of dynamics penalties (see Table 7), based on the assumption that the remaining performance gap compared to online RL is solely due to inaccuracies in the learned dynamics model.
+This understanding naturally implies that improving the dynamics model should also improve performance.
+Surprisingly, we find that existing offline model-based methods completely fail if the learned dynamics model is replaced with the true, error-free dynamics model, while keeping everything else the same (see Figure 1).
+Under the true dynamics, the only difference to online RL is that in online RL, data is sampled as full-length episodes, while in offline model-based RL, data is instead sampled as $k$-step rollouts, starting from a state in the original offline dataset, with rollout length $k$ limited to avoid accumulating model errors.
+Failure under the true model therefore highlights that truncating rollouts has critical and previously-overlooked consequences.
+We find that this rollout truncation leads to a set of states which, under any policy, can only be reached in the final rollout step (see red in Figure 2).
+The existence of these edge-of-reach states is problematic as it means Bellman updates (see Equation 1) use target values that are never themselves trained.
+This is illustrated in Figure 2, and described in detail in Section 3.
+This effective “bootstrapping from the void” triggers a catastrophic breakdown in $Q$-learning.
+Concisely, this issue can be viewed as all actions from edge-of-reach states remaining out-of-reach over training.
+Hence, contrary to common understanding, the out-of-sample action problem central to model-free methods is not fully resolved by a model-based approach.
+In fact, in Section 3.4 we provide detailed analysis suggesting that this is the predominant source of issues on the standard D4RL benchmark.
+In Section 6.5 we consequently reexamine how existing methods work and find they are indirectly and unintentionally addressing this issue (rather than model errors as claimed).
+We have the following problem: Indirectly addressing the edge-of-reach problem means existing model-based methods have a fragile and unforeseen dependence on model quality and fail catastrophically as dynamics models improve.
+Since model improvements are inevitable, and since tuning of hyperparameters is particularly impractical in offline RL, this is a substantial practical barrier.
+In light of this, we propose Reach-Aware Value Learning (RAVL), a simple and robust method that directly addresses the edge-of-reach problem.
+RAVL achieves strong performance on standard benchmarks, and moreover continues to perform even with error-free, uncertainty-free dynamics models (Section 6).
+Our model-based method bears close connections to model-free approaches, and hence in Section A.2 we present a unified perspective of these two previously disjoint subfields.
+Algorithm 1: Pseudocode for the base procedure used in offline model-based methods.
+In summary: (1) Train a dynamics model $\widehat{M}$ on $\mathcal{D}_{\textrm{offline}}$, then (2) Train an agent (often SAC) with $k$-step rollouts in the learned model starting from $s\in\mathcal{D}_{\textrm{offline}}$.
+Existing methods consider issues to be due to errors in $\widehat{M}$ and hence introduce dynamics uncertainty penalties (see Table 7).
+We find the predominant source of issues to be the edge-of-reach problem, and hence instead propose using value pessimism (RAVL) (see Section 5).
+Algorithm 1 Base model-based algorithm (MBPO) + Additions in existing methods and RAVL (ours)
+1:  Require: Offline dataset $\mathcal{D}_{\textrm{offline}}$
+2:  Require: Dynamics model $\widehat{M}$ = ($\widehat{T}$, $\widehat{R}$) trained on $\mathcal{D}_{\textrm{offline}}$. Augment $\bm{\widehat{M}}$ with an uncertainty penalty *
+3:  Specify: Rollout length $k\geq 1$, real data ratio $r\in[0,1]$
+4:  Initialize: Replay buffer $\mathcal{D}_{\textrm{rollouts}}=\emptyset$, policy $\pi_{\theta}$, value function $Q_{\phi}$ (both from random)
+5:  for $\textrm{epochs}=1,\dots$ do
+6:     (Collect data) Starting from states in $\mathcal{D}_{\textrm{offline}}$, collect $k$-step rollouts in $\widehat{M}$ with $\pi_{\theta}$. Store data in $\mathcal{D}_{\textrm{rollouts}}$
+7:     (Train agent) Train $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\textrm{rollouts}}\cup\mathcal{D}_{\textrm{offline}}$ (mixed with ratio $r$) Add $\mathbf{Q}$-value pessimism
+8:  end for  *This uncertainty penalty collapses to zero with the true error-free dynamics ($\bm{\widehat{M}}=\bm{M}$) (see experiments Table 2).
 2 Background
 Reinforcement Learning.
-We consider the standard RL framework (Sutton & Barto, 2018), in which the environment is formulated as a Markov Decision Process, $M=(\mathcal{S},\mathcal{A},T,R,\mu_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, $T(s^{\prime}|s,a)$ and $R(s,a)$ denote the transition and reward dynamics, $\mu_{0}$ the initial state distribution, and $\gamma\in(0,1)$ is the discount factor.
+We consider the standard RL framework [31], in which the environment is formulated as a Markov Decision Process, $M=(\mathcal{S},\mathcal{A},T,R,\mu_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, $T(s^{\prime}|s,a)$ and $R(s,a)$ denote the transition and reward dynamics, $\mu_{0}$ the initial state distribution, and $\gamma\in(0,1)$ is the discount factor.
 The goal in reinforcement learning is to learn a policy $\pi(a|s)$ that maximizes the expected discounted return $\mathbb{E}_{\mu_{0},\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$.
 Actor-Critic Algorithms.
-The broad class of algorithms we consider are actor-critic (Konda & Tsitsiklis, 1999) methods which jointly optimize a policy $\pi$ and state-action value function ($Q$-function).
-Actor-critic algorithms iterate between fitting the $Q$-function to the current policy (policy evaluation) and improving the policy relative to the $Q$-function (policy improvement).
-The policy evaluation step relies on approximate dynamic programming with Bellman updates, updating the $Q$-function using some dataset $\mathcal{D}$ as follows:
+The broad class of algorithms we consider are actor-critic [15] methods which jointly optimize a policy $\pi$ and state-action value function ($Q$-function).
+Given a dataset $\mathcal{D}$ of (state, action, reward, nextstate) transitions, actor-critic algorithms iterate between two steps:
+1.
+(Policy evaluation) Update $Q_{\phi}$ to fit the current policy using the following Bellman update: $\begin{split}\text{{\color[rgb]{0.8,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.0}\footnotesize{Key expression:}}}\qquad\underbrace{Q_{\phi}(s,a)}_{\textbf{\tiny{\color[rgb]{0.8,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.0}Update value at $\bm{(s,a)}$}}}\leftarrow\underbrace{r+\gamma Q_{\phi}(s^{\prime},\pi_{\theta}(s^{\prime}))}_{\textbf{\tiny{\color[rgb]{0.8,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.0}Using value at $\bm{(s^{\prime},\pi_{\theta}(s^{\prime}))}$}}}\qquad(s,a,r,s^{\prime})\sim\mathcal{D}\end{split}$ (1)
+2.
+(Policy improvement) Update $\pi_{\theta}$ to increase the current $Q$-value predictions: $Q_{\phi}(s,\pi_{\theta}(s))$ [^3] This step contains an implicit $\max$ operation which we refer to in Section 3.2..
+Offline RL and the Out-of-Sample Action Problem.
+In offline RL, training must rely on only a fixed dataset of transitions $\mathcal{D}_{\textrm{offline}}=\{(s,a,r,s^{\prime})^{(i)}\}_{i=1,\dots,N}$ collected by some policy $\pi^{\beta}$.
+The central problem in this offline setting is the out-of-sample [^4] We use the terms ‘out-of-sample’ and ‘out-of-distribution’ interchangeably, as is done in the literature.
+action problem: The values of $Q_{\phi}$ being updated are at any state-actions that appear as $(s,a)$ in $\mathcal{D}_{\textrm{offline}}$.
+However, the targets of these updates rely on values at $(s^{\prime},a^{\prime})$ where $a^{\prime}\sim\pi_{\theta}$ (see Equation 1).
+Since $a^{\prime}\sim\pi_{\theta}$ (on-policy) whereas $a\sim\pi^{\beta}$ (off-policy), $(s^{\prime},a^{\prime})$ may not appear in $\mathcal{D}_{\textrm{offline}}$, and hence may itself have never been updated.
+This means that updates can involve bootstrapping from ‘out-of-sample’ and hence arbitrarily misestimated values and thus lead to misestimation being propagated over the entire state-action space.
+The $\max$ operation (implicit in the policy improvement step) further acts to exploit any misestimation, converting misestimation into extreme pathological overestimation.
+As a result $Q$-values often tend to increase unboundedly over training, while performance collapses entirely [23, 18].
+Model-Based Offline RL.
+Model-based methods [30] aim to solve the out-of-sample issue by allowing the agent to collect additional synthetic data in a learned dynamics model.
+They generally share the same base procedure as described in Algorithm 1.
+This involves first training an approximate dynamics model $\widehat{M}=(\mathcal{S},\mathcal{A},\widehat{T},\widehat{R},\mu_{0},\gamma)$ on $\mathcal{D}_{\textrm{offline}}$.
+Here, $\widehat{T}(s^{\prime}|s,a)$ and $\widehat{R}(s,a)$ denote the learned transition and reward functions, commonly realized as a deep ensemble [5, 22].
+Following this an agent is trained using an online RL algorithm (typically SAC [11]), for which data is sampled as $k$-step trajectories (termed rollouts) under the current policy, starting from states in the offline dataset $\mathcal{D}_{\textrm{offline}}$.
+The base procedure of training a SAC agent with model rollouts does not work out of the box.
+Existing methods attribute this as due to dynamics model errors.
+Consequently, a broad class of methods propose augmenting the learned dynamics model $\widehat{M}$ with some form of dynamics uncertainty penalty, often based on variance over the ensemble dynamics model [36, 29, 14, 24, 2].
+We present some explicit examples in Table 7.
+Crucially, all of these methods assume that, under the true error-free model, no intervention should be needed, and hence all the uncertainty penalties collapse to zero.
+In the following section, we show that this assumption leads to catastrophic failure.
+Figure 2: The previously unnoticed edge-of-reach problem.
+Left illustrates the base procedure used in offline model-based RL, whereby synthetic data is sampled as $k$-step trajectories “rollouts” starting from a state in the original offline dataset.
+Edge-of-reach states are those that can be reached in $k$-steps, but which cannot (under any policy) be reached in less than $k$-steps.
+We depict the data collected with two rollouts, one ending in $s_{k}=D$, and the other with $s_{k}=C$.
+Right then shows this data arranged into a dataset of transitions as used in $Q$-updates.
+State $D$ is edge-of-reach and hence appears in the dataset as $s^{\prime}$ but never as $s$.
+Bellman updates therefore bootstrap from $D$, but never update the value at $D$ (see Equation 1).
+(For comparison consider state $C$: $C$ is also sampled at $s_{k}$, but unlike $D$ it is not edge-of-reach, and hence is also sampled at $s_{i<k}$ meaning it is updated and hence does not cause issues.)
+3 The Edge-of-Reach Problem
+In the following section, we formally introduce the edge-of-reach problem.
+We begin by showing the empirical failure of SOTA offline model-based methods on the true environment dynamics.
+Next, we present our edge-of-reach hypothesis for this, including intuition, empirical evidence on the main D4RL benchmark [7], and theoretical proof of its effect on offline model-based training.
+3.1 Surprising Failure with the True Dynamics
+As described in Sections 1 and 2, prior works view offline model-based RL as online RL in an approximate dynamics model.
+Based on this understanding they propose various forms of dynamics uncertainty penalties to address model errors (see Table 7).
+This approach is described simply as: “two steps: (a) learning a pessimistic MDP (P-MDP) using the offline dataset; (b) learning a near-optimal policy in this P-MDP”[14].
+This shared base procedure is shown in Algorithm 1, where “P-MDP” refers to the learned dynamics model with a penalty added to address model errors.
 […]
-Offline RL and the Out-of-Sample Problem.
-In online RL, the agent is able to collect new on-policy data throughout training.
-By contrast, in the offline setting, the algorithm has access only to a fixed offline dataset consisting of transition tuples $\mathcal{D}_{\textrm{offline}}=\{(s,a,r,s^{\prime})^{(i)}\}_{i=1,\dots,N}$ collected by some behavioral policies $\pi^{\beta}$.
-The central problem in offline learning is the out-of-sample [^3] We use the terms ‘out-of-sample’ and ‘out-of-distribution’ interchangeably, as is done in the literature.
-action problem: Target actions being sampled on-policy $a^{\prime}\sim\pi^{j}$ means state-actions $(s^{\prime},a^{\prime})$ may not appear as $(s,a)$ in the dataset (see Equation 1).
-Values at these out-of-sample state-actions are therefore never themselves updated and hence likely to be arbitrarily misestimated, but yet they are used for the targets in Bellman updates.
-Updates therefore propagate this misestimation to previous (in-sample) state-actions leading to a complete breakdown in $Q$-learning.
-Further, the $\max$ operation implicit in the policy improvement serves to pick out the highest misestimated $Q$-values such that $Q$-values become extremely overestimated, often increasing unboundedly over training (Levine et al., 2020; Kumar et al., 2019).
-Offline Model-Based RL Model-based methods (Sutton, 1991) aim to solve the out-of-sample issue by allowing the agent to collect additional synthetic data in a learned dynamics model $\widehat{M}=(\mathcal{S},\mathcal{A},\widehat{T},\widehat{R},\mu_{0},\gamma)$.
-$\widehat{T}(s^{\prime}|s,a)$ and $\widehat{R}(s,a)$ denote the approximate transition and reward functions, and are commonly realized as a deep ensemble (Chua et al., 2018; Lakshminarayanan et al., 2017).
-Methods typically use the MBPO procedure (Janner et al., 2019) shown in Algorithm 1, with synthetic data sampled as short $k$-step trajectories (termed rollouts) under the current policy, starting from states in the offline dataset $\mathcal{D}_{\textrm{offline}}$.
-Algorithm 1: Pseudocode for the base procedure used in offline model-based methods (MBPO, Janner et al. (2019)).
-In summary:
+3.2 The Edge-Of-Reach Hypothesis (Illustrated in Figure 2)
+Failure under the error-free dynamics reveals that “model errors” cannot explain all the problems in offline model-based RL.
+Instead, it highlights that there must be a second issue.
+On investigation we find this to be the “edge-of-reach problem.” We begin with the main intuition: in offline model-based RL, synthetic data $\mathcal{D}_{\textrm{rollouts}}$ is generated as short $k$-step rollouts starting from states in the original dataset $\mathcal{D}_{\textrm{offline}}$ (see Algorithm 1).
+Crucially, Figure 2 (left) illustrates how, under this procedure, there can exist some states which can be reached in the final rollout step, but which cannot - under any policy - be reached earlier.
+These edge-of-reach states triggers a breakdown in learning, since even with the ability to collect unlimited data, the agent is never able to reach these states ‘in time’ to try actions from them, and hence is free to ‘believe that these edge-of-reach states are great.’
+More concretely: Figure 2 (right) illustrates how being sampled only at the final rollout step means edge-of-reach states will appear in the resulting dataset $\mathcal{D}_{\textrm{rollouts}}$ as nextstates $s^{\prime}$, but will never appear in the dataset as states $s$.
+Crucially, in Equation 1, we see that values at states $s$ are updated, while values at $s^{\prime}$ are used for the targets of these updates.
+Edge-of-reach states are therefore used for targets, but are never themselves updated, meaning that their values can be arbitrarily misestimated.
+Updates consequently propagate misestimation over the entire state-action space.
+Furthermore, the $\max$ operation in the policy improvement step (see Section 2) exploits any misestimation by picking out the most heavily overestimated values [23].
+The misestimation is therefore turned into overestimation, resulting in the value explosion seen in Figure 3.
+Thus, contrary to common understanding, the out-of-sample action problem key in model-free offline RL can be seen to persist in model-based RL.
+Edge-of-Reach Hypothesis: Limited horizon rollouts from a fixed dataset leads to ‘edge-of-reach’ states: states which are used as targets for Bellman-based updates, but which are never themselves updated.
+The resulting maximization over misestimated values in Bellman updates causes pathological value overestimation and a breakdown in $Q$-learning.
+When is this an issue in practice?
+Failure via this mode requires (a) the existence of edge-of-reach states, and (b) such edge-of-reach states to be sampled.
+The typical combination of $k\ll H$ along with a limited pool of starting states ($s\in\mathcal{D}_{\textrm{offline}}$) means the rollout distribution is unlikely to sufficiently cover the full state space $\mathcal{S}$, thus making (a) likely.
+Moreover, we observe pathological ‘edge-of-reach seeking’ behavior in which the agent appears to ‘seek out’ edge-of-reach states due to this being the source of overestimation, thus making (b) likely.
+This behaviour is discussed further in Section 4.
+In Appendix A, we give a thorough and unified view of model-free and model-based offline RL, dividing the problem into independent conditions for states and actions and examining when we can expect the edge-of-reach problem to be significant.
+3.3 Definitions and Formalization
+Definition 1 (Edge-of-reach states).
+Consider a deterministic transition model $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, rollout length $k$, and some distribution over starting states $\nu_{0}$.
+For some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$, rollouts are then generated according to $s_{0}\sim\nu_{0}(\cdot)$, $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim T(\cdot|s_{t},a_{t})$ for $t=0,\dots,k-1$, giving $(s_{0},a_{0},s_{1},\dots,s_{k})$.
+Let us use $\rho_{t,\pi}(s)$ to denote the marginal distributions over $s_{t}$.
+We define a state $s\in S$ edge-of-reach with respect to ($T$, $k$, $\nu_{0}$) if: for $t=k$, $\exists\>\pi$ s.t.
+$\rho_{t,\pi}(s)>0$, but, for $t=1,\dots,k-1$ and $\forall\>\pi$, $\rho_{t,\pi}(s)=0$.
+In our case, $\nu_{0}$ is the distribution of states in $\mathcal{D}_{\textrm{offline}}$.
+In Appendix B, we include an extension to stochastic transition models, proof of how errors can consequently propagate to all states, and a discussion of the practical implications.
+3.4 Empirical evidence on the D4RL benchmark
+There are two potential issues in offline model-based RL: (1) dynamics model errors and subsequent model exploitation, and (2) the edge-of-reach problem and subsequent pathological value overestimation.
+We ask: Which is the true source of the issues observed in practice?
+While (1) is stated as the sole issue and hence the motivation in prior methods, the “failure” results presented in Figure 1 and Table 1 are strongly at odds with this explanation.
+Furthermore, in Table 8, we examine the model rewards sampled by the agent over training.
+For (1) to explain the $Q$-values seen in Figure 3, the sampled rewards in Table 8 would need to be on the order of $10^{8}$.
+Instead, however, they remain less than $10$, and are not larger than the true rewards, meaning (1) does not explain the observed value explosion (see Section D.2).
+By contrast, the edge-of-reach-induced pathological overestimation mechanism of (2) exactly predicts this value explosion, with the observations very closely resembling those of the analogous model-free out-of-sample problem [18].
+Furthermore, (2) is consistent with the “failure” observations in Table 1.
+Finally, we again highlight the discussion in Section 3.2 where we explain why the edge-of-reach problem can be expected to occur in practice.
+Figure 3: The base procedure results in poor performance (left) with exponential increase in $Q$-values (right) on the D4RL benchmark.
+Approx $Q^{*}$ indicates the $Q$-value for a normalized score of 100 (with $\gamma=0.99$).
+Results are shown for Walker2d-medexp (6 seeds), but we note similar trends across other D4RL datasets.
+4 Analysis with a Simple Environment
+In the previous section, we presented the edge-of-reach hypothesis as an explanation of why existing methods fail under the true dynamics.
+In this section, we construct a simple environment to empirically confirm this hypothesis.
+We first reproduce the observation seen in Figures 1 and 3 and Table 1 of the base offline model-based procedure resulting in exploding $Q$-values and failure to learn despite using the true dynamics model.
+Next, we verify that edge-of-reach states are the source of this problem by showing that correcting value estimates only at these states is sufficient to resolve the issues.
+Figure 4: Experiments on the simple environment, illustrating the edge-of-reach problem and potential solutions.
+(a) Reward function, (b) final (failed) policy with naïve application of the base procedure (see Algorithm 1), (c) final (successful) policy with patching in oracle $Q$-values for edge-of-reach states, (d) final (successful) policy with RAVL, (e) returns evaluated over training, (f) mean $Q$-values evaluated over training.
+4.1 Setup
+[…]
+In applying the offline model-based procedure (see Algorithm 1) to this (true) environment we have precisely the same setup as in Section 3.1, where again the only difference compared to online RL is the use of truncated $k$-step rollouts with $k=10$ (compared to full horizon $H=30$).
+This small change results in the existence of edge-of-reach states (those between the red and orange boxes).
+4.2 Observing Pathological Value Overestimation
+Exactly as with the benchmark experiments in Section 3.1, the base model-based procedure fails despite using the true dynamics (see blue Figure 4, Base), and the $Q$-values grow unboundedly over training (compare Figure 4f and Figure 3b).
+Looking at the rollouts sampled over training (see Figure 7) we see the following behavior:
+(Before 25 epochs) Performance initially increases.
+(Between 25 and 160 epochs) Value misestimation takes over, and the policy begins to aim toward unobserved state-actions (since their values can be misestimated and hence overestimated).
+(After 160 epochs) This ‘edge-of-reach seeking’ behavior compounds with each epoch, leading the agent to eventually reach edge-of-reach states.
+From this point onwards, the agent samples edge-of-reach states at which it never receives any corrective feedback.
+The consequent pathological value overestimation results in a complete collapse in performance.
+In Figure 4b we visualize the final policy and see that it completely ignores the reward function, aiming instead towards an arbitrary edge-of-reach state.
+4.3 Verifying the Hypothesis Using Value Patching
+Our hypothesis is that the source of this failure is value misestimation at edge-of-reach states.
+Our Base-OraclePatch experiments (see yellow Figure 4) verify this by showing that patching in the correct values solely at edge-of-reach states is sufficient to completely solve the problem.
+This is particularly compelling as in practice we only corrected values at 0.4% of states over training.
+In Section 5 we introduce our practical method RAVL, which Figures 4 green and 7 show has an extremely similar effect to that of the ideal but practically impossible Base-OraclePatch intervention.
+5 RAVL: Reach-Aware Value Learning
+As verified in Section 4.3, issues stem from value overestimation at edge-of-reach states. To resolve this, we therefore need to (A) detect and (B) prevent overestimation at edge-of-reach states.
+(A) Detecting edge-of-reach states: As illustrated in Figure 2 (right), edge-of-reach states are states at which the $Q$-values are never updated, i.e., those that are out-of-distribution (OOD) with respect to the training distribution of the $Q$-function, $s\in\mathcal{D}_{\textrm{rollouts}}$.
+A natural solution for OOD detection is measuring high variance over a deep ensemble [22].
+We can therefore detect edge-of-reach states using an ensemble of $Q$-functions.
+We demonstrate that this is effective in Figure 5.
+(B) Preventing overestimation at these states: Once we have detected edge-of-reach states, we may simply apply value pessimism methods from the offline model-free literature.
+Our choice of an ensemble for part (A) conveniently allows us to minimize over an ensemble of $Q$-functions, which effectively adds value pessimism based on ensemble variance.
+Our resulting proposal is Reach-Aware Value Learning (RAVL).
+Concretely, we take the standard offline model-based RL procedure (see Algorithm 1), and simply exchange the dynamics pessimism penalty for value pessimism using minimization over an ensemble of $N$ $Q$-functions:
+$\begin{split}\qquad\qquad Q_{\phi}^{n}(s,a)\leftarrow r+\gamma\min_{i=1,\dots,N}Q_{\phi}^{i}(s^{\prime},\pi_{\theta}(s^{\prime}))\quad\text{ for }n=1,\dots,N\end{split}$ (2)
+We include EDAC’s [1] ensemble diversity regularizer, and in Section 7 we discuss how the impact of value pessimism differs significantly in the model-based (RAVL) vs model-free (EDAC) settings.
+6 Empirical Evaluation
+In this section, we begin by analyzing RAVL on the simple environment from Section 4.
+Next, we look at the standard D4RL benchmark, first confirming that RAVL solves the failure seen with the true dynamics, before then demonstrating that RAVL achieves strong performance with the learned dynamics.
+In Appendix F, we include additional results on the challenging pixel-based V-D4RL benchmark on which RAVL now represents a new state-of-the-art.
+Finally in Section 6.5 we reexamine prior model-based methods and explain why they may work despite not explicitly addressing the edge-of-reach problem.
+We provide full hyperparameters in Section C.2 and ablations showing that RAVL is stable over hyperparameter choice in Section G.2.
+Figure 5: RAVL’s effective penalty of $Q$-ensemble variance on the environment in Section 4, showing that - as intended - edge-of-reach states have significantly higher penalty than within-reach states.
+6.1 Simple Environment
+Testing on the simple environment from Section 4 (see green Figure 4), we observe that RAVL behaves the same as the theoretically optimal but practically impossible Base-OraclePatch method.
+Moreover, in Figure 5, we see that the $Q$-value variance over the ensemble is significantly higher for edge-of-reach states, meaning RAVL is detecting and penalizing edge-of-reach states exactly as intended.
+6.2 D4RL with the True Dynamics
+Next, we demonstrate that RAVL works without dynamics uncertainty and solves the ‘failure’ observed in Section 3.1.
+Table 1 shows results on the standard offline benchmark D4RL [7] MuJoCo [34] v2 datasets with the true (zero error, zero uncertainty) dynamics.
+We see that RAVL learns the near-optimal policies, while existing methods using the base model-based procedure (Algorithm 1) completely fail.
+In Section 6.5 we examine how this is because existing methods overlook the critical edge-of-reach problem and instead only accidentally address it using dynamics uncertainty metrics.
+In the absence of model uncertainty, these methods have no correction for edge-of-reach states and hence fail dramatically.
+Table 1: True dynamics (zero error, zero uncertainty) Existing model-based methods are presented as different approaches for dealing with dynamics model errors.
+Surprisingly, however, all existing methods fail in the absence of dynamics errors (when the learned approximate model is replaced by the true model).
+This reveals that existing methods are unintentionally using their dynamics uncertainty estimates to address the previously unnoticed edge-of-reach problem.
+By contrast, RAVL directly addresses the edge-of-reach problem and hence does not fail when dynamics uncertainty is zero.
+Experiments are on the D4RL MuJoCo v2 datasets.
+Statistical significance highlighted (6 seeds).
+**Note that while labeled as ‘MOBILE’, the results with the true dynamics will be identical for any other dynamics penalty-based method since penalties under the true model are all zero (see Table 7).
+Hence these results indicate the failure of all existing dynamics uncertainty-based methods.
+MOBILE (SOTA)** RAVL
+Dynamics model Approximate True Approximate True
+Halfcheetah medium 74.6$\pm$1.2 72.2$\pm$4.1 ($\downarrow$7%) 78.7$\pm$2.0 76.4$\pm$3.4
+mixed 71.7$\pm$1.2 72.2$\pm$4.1 ($\uparrow$1%) 74.9$\pm$2.0 77.3$\pm$2.2
+medexp 108.2$\pm$2.5 84.9$\pm$17.6 ($\downarrow$22%) 102.1$\pm$8.0 109.3$\pm$1.3
+Hopper medium 106.6$\pm$0.6 17.2$\pm$11.1 ($\downarrow$84%) 90.6$\pm$11.9 101.6$\pm$0.9
+mixed 103.9$\pm$1.0 71.5$\pm$32.2 ($\downarrow$31%) 103.1$\pm$1.1 101.1$\pm$0.5
+medexp 112.6$\pm$0.2 6.1$\pm$6.8 ($\downarrow$96%) 110.1$\pm$1.8 107.2$\pm$0.9
+Walker2d medium 87.7$\pm$1.1 7.2$\pm$1.6 ($\downarrow$92%) 86.3$\pm$1.6 101.2$\pm$2.3
+mixed 89.9$\pm$1.5 7.9$\pm$1.9 ($\downarrow$91%) 83.0$\pm$2.5 86.4$\pm$3.0
+medexp 115.2$\pm$0.7 7.7$\pm$2.1 ($\downarrow$93%) 115.5$\pm$2.4 103.5$\pm$1.2
+6.3 D4RL with Learned Dynamics
+Next, we show that RAVL also performs well with a learned dynamics model.
+Figure 3 shows RAVL successfully stabilizes the $Q$-value explosion of the base procedure, and Table 2 shows RAVL largely matches SOTA, while having significantly lower runtime (see Section 6.4).
+RAVL gives much higher performance on the Halfcheetah mixed and medium datasets than its model-free counterpart EDAC, and in Section 6.5, we discuss why we would expect the effect of value pessimism in the model-based setting (RAVL) to inherently offer much more flexibility than in the model-free setting (EDAC).
+Table 2: A comprehensive evaluation of RAVL over the standard D4RL MuJoCo benchmark.
+We show the mean and standard deviation of the final performance averaged over 6 seeds.
+Our simple approach largely matches the state-of-the-art without any explicit dynamics penalization and hence works even in the absence of model uncertainty (where dynamics uncertainty-based methods fail) (see Table 1).
+Model-Free Model-Based
+Environment BC CQL EDAC MOPO COMBO RAMBO MOBILE RAVL (Ours)
+Halfcheetah random 2.2 31.3 28.4 38.5 38.8 39.5 39.3 34.4$\pm$2.0
+medium 43.2 46.9 65.9 73.0 54.2 77.9 74.6 78.7$\pm$2.0
+mixed 37.6 45.3 61.3 72.1 55.1 68.7 71.7 74.9$\pm$2.0
+medexp 44.0 95.0 106.3 90.8 90.0 95.4 108.2 102.1$\pm$8.0
+Hopper random 3.7 5.3 25.3 31.7 17.9 25.4 31.9 31.4$\pm$0.1
+medium 54.1 61.9 101.6 62.8 97.2 87.0 106.6 90.6$\pm$11.9
+mixed 16.6 86.3 101.0 103.5 89.5 99.5 103.9 103.1$\pm$1.1
+medexp 53.9 96.9 110.7 81.6 111.1 88.2 112.6 110.1$\pm$1.8
+Walker2d random 1.3 5.4 16.6 7.4 7.0 0.0 17.9 17.1$\pm$7.9
+medium 70.9 79.5 92.5 84.1 81.9 84.9 87.7 86.3$\pm$1.6
+mixed 20.3 76.8 87.1 85.6 56.0 89.2 89.9 83.0$\pm$2.5
+medexp 90.1 109.1 114.7 112.9 103.3 56.7 115.2 115.5$\pm$2.4
+We additionally include results for the challenging pixel-based V-D4RL benchmark for which latent-space models are used (in Appendix F), and accompanying ablation experiments (in Section G.2).
+In this setting, RAVL represents a new SOTA, and the results are particularly notable as the pixel-based setting means the base algorithm (DreamerV2) uses model rollouts in an imagined latent space (rather than the original state-action space as in MBPO).
+The results therefore give promising evidence that RAVL is able to generalize well to different representation spaces.
+6.4 Runtime Discussion
+Vectorized ensembles can be scaled with extremely minimal effect on the runtime (see Table 10).
+This means that, per epoch, RAVL is approximately 13% faster than the SOTA (MOBILE, due to MOBILE needing multiple extra forward passes to compute its uncertainty penalty).
+Furthermore, in total, we find that RAVL reliably requires 3×3$\times$3 × fewer epochs to converge on all but the medexp datasets, meaning the total runtime is approximately 70% faster than SOTA (see Section G.1).
+6.5 How can prior methods work despite overlooking the edge-of-reach problem?
+While ostensibly to address model errors, we find that existing dynamics penalties accidentally address the edge-of-reach problem: In Figure 6 we see a positive correlation between the penalties used in dynamics uncertainty methods and RAVL’s effective penalty of value ensemble variance.
+This may be expected, as dynamics uncertainty will naturally be higher further away from $\mathcal{D}_{\textrm{offline}}$, which is also where edge-of-reach states are more likely to lie.
+In Section 3.4 we present evidence suggesting that the edge-of-reach problem is likely the dominant source of issues on the main D4RL benchmark, thus indicating that the dynamics uncertainty penalties of existing methods are likely indirectly addressing the edge-of-reach problem.
+In general, dynamics errors are a second orthogonal problem and RAVL can be easily combined with appropriate dynamics uncertainty penalization [36, 29, 24] for environments where this is a significant issue.
+7 Related Work
+Model-Based Methods.
+Existing offline model-based methods present dynamics model errors and consequent model exploitation as the sole source of issues.
+A broad class of methods therefore propose reward penalties based on the estimated level of model uncertainty [36, 14, 29, 24, 2], typically using variance over a dynamics ensemble.
+Rigter et al.
+[27] aim to avoid model exploitation by setting up an adversarial two-player game between the policy and model.
+Finally, most related to our method, COMBO [37], penalizes value estimates for state-actions outside model rollouts.
+However, similarly to Yu et al.
+[36], COMBO is theoretically motivated by the assumption of infinite horizon model rollouts, which we show overlooks serious implications.
+Critically, in contrast to our approach, none of these methods address the edge-of-reach problem and thus they fail as environment models become more accurate.
+A related phenomenon of overestimation stemming from hallucinated states has been observed in online model-based RL[12].
+Offline Model-Free Methods.
+Model-free methods can broadly be divided into two approaches to solving the out-of-sample action problem central to offline model-free RL (see Section 2): action constraint methods and value pessimism-based methods.
+Action constraint methods [17, 18, 8] aim to avoid using out-of-sample actions in the Bellman update by ensuring selected actions are close to the dataset behavior policy $\pi^{\beta}$.
+By contrast, value pessimism-based methods aim to directly regularize the value function to produce low-value estimates for out-of-sample state-actions [19, 16, 1].
+The edge-of-reach problem is the model-based equivalent of the out-of-sample action problem, and this unified understanding allows us to transfer ideas directly from model-free literature.
+RAVL is based on EDAC’s [1] use of minimization over a $Q$-ensemble [35, 9], and applies this to model-based offline RL.
+What is the effect of value pessimism in the model-based vs model-free settings?
+EDAC [1] can be seen as RAVL’s model-free counterpart, however, the impact of value pessimism in the model-free vs the model-based settings is notably different.
+Recall that, with the ensemble $Q$-function trained on $(s,a,r,s\textquoteright)\sim\mathcal{D}_{\Box}$, the state-actions that are penalized (due to being outside the training distribution) are any $(s\textquoteright,a\textquoteright)$ ’ ) that are out-of-distribution with respect to the $(s,a)$’s in the dataset $\mathcal{D}_{\Box}$.
+Recall also that updates use values at $(s^{\prime},a^{\prime})$ where $a^{\prime}\sim\pi_{\theta}$ (see Equation 1).
+In the model-free case (EDAC): The dataset is $\mathcal{D}_{\Box}=\mathcal{D}_{offline}$, meaning the actions $a$ are effectively sampled from the dataset behavior policy (off-policy), whereas the actions $a\textquoteright$ ’ are sampled on-policy.
+This means that EDAC penalizes any $(s\textquoteright,a\textquoteright)$ ’ ) where $a\textquoteright$ ’ differs significantly from the behavior policy.
+In the model-based case (RAVL): The dataset is $\mathcal{D}_{\Box}=\mathcal{D}_{rollouts}$, meaning now both $a$ and $a\textquoteright$ ’ are sampled on-policy.
+As a result, the only $(s\textquoteright,a\textquoteright)$ ’ ) which will now be out-of-distribution with respect to $(s,a)$ (and hence penalized) are those where the state $s\textquoteright$ ’ is out-of-distribution with respect to the states $s$ in $\mathcal{D}_{rollouts}$.
+This happens when $s\textquoteright$ ’ is reachable only in the final step of rollouts, i.e.
+exactly when $s\textquoteright$ ’ is “edge-of-reach” (as illustrated in Figure 2).
+Compared to EDAC, RAVL can therefore be viewed as “relaxing” the penalty and giving the agent freedom to learn a policy that differs significantly from the dataset behavior policy.
+This distinction is covered in detail in Appendix A.
+8 Conclusion
+This paper investigates how offline model-based methods perform as dynamics models become more accurate.
+As an interesting hypothetical extreme, we test existing methods with the true error-free dynamics.
+Surprisingly, we find that all existing methods fail.
+This reveals that using truncated rollout horizons (as per the shared base procedure) has critical and previously overlooked consequences stemming from the consequent existence of ‘edge-of-reach’ states.
+We show that existing methods are indirectly and accidentally addressing this edge-of-reach problem (rather than addressing model errors as stated), and hence explain why they fail catastrophically with the true dynamics.
+This problem reveals close connections between model-based and model-free approaches and leads us to present a unified perspective for offline RL.
+Based on this, we propose RAVL, a simple and robust method that achieves strong performance across both proprioceptive and pixel-based benchmarks.
+Moreover, RAVL directly addresses the edge-of-reach problem, meaning that - unlike existing methods - RAVL does not fail under the true environment model, and has the practical benefit of not requiring dynamics uncertainty estimates.
+Since improvements to dynamics models are inevitable, we believe that resolving the brittle and unanticipated failure of existing methods under dynamics model improvements is an important step towards ‘future-proofing’ offline RL.
+9 Limitations
+Since dynamics models for the main offline RL benchmarks are highly accurate, the edge-of-reach effects dominate, and RAVL is sufficient to stabilize model-based training effectively without any explicit dynamics uncertainty penalty.
+In general, however, edge-of-reach issues could be mixed with dynamics error, and understanding how to balance these two concerns would be useful future work.
+Further, we believe that studying the impact of the edge-of-reach effect in a wider setting could be an exciting direction, for example investigating its effect as an implicit exploration bias in online RL.
+Acknowledgments and Disclosure of Funding
+AS is supported by the EPSRC Centre for Doctoral Training in Modern Statistics and Statistical Machine Learning (EP/S023151/1).
+The authors would like to thank the conference reviewers, Philip J.
+Ball, and Shimon Whiteson for their helpful feedback and discussion.
+Supplementary Material
 […]
 […]
-Existing methods consider issues to be due to model errors and hence introduce dynamics uncertainty penalties.
-We instead reveal that the true source of issues is the edge-of-reach problem, and propose a value pessimism approach to directly address this.
-Algorithm 1 Base model-based algorithm (MBPO)
-+ Additions in existing methods and RAVL (ours)
-1:  Require:Offline dataset $\mathcal{D}_{\textrm{offline}}$earned dynamics model $\widehat{M}$ = ($\widehat{T}$, $\widehat{R}$) trained on $\mathcal{D}_{\textrm{offline}}$ynamics uncertainty penalty added to $\widehat{M}$
-2:  Specify:Rollout length $k\geq 1$, real data ratio $r\in[0,1]$
-3:  Initialize:Replay buffer $\mathcal{D}_{\textrm{rollouts}}=\emptyset$, policy $\pi_{\theta}$, value function $Q_{\phi}$
-4:  for $\textrm{epochs}=1,\dots$ do
-5:     (Collect data)Starting from states in $\mathcal{D}_{\textrm{offline}}$, collect $k$-step rollouts in $\widehat{M}$ with $\pi_{\theta}$. Store data in $\mathcal{D}_{\textrm{rollouts}}$
-6:     (Train agent)Train $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\textrm{rollouts}}\cup\mathcal{D}_{\textrm{offline}}$ mixed with ratio $r$ with value pessimism applied to the $Q$-function
-7:  end for*Vanishes with an Oracle model (Table 1) as $\bm{\widehat{M}}$ replaced by true $\bm{M}$
-3 Related Work
-Overestimation bias in $Q$-learning.
-$Q$-learning is a popular basis for many off-policy RL algorithms, however, it has been shown to suffer from an overestimation bias even in the online case (van Hasselt et al., 2016).
-In online RL, Fujimoto et al.
-(2018) demonstrate taking the minimum over two independently trained $Q$-functions to be an effective solution, and extending this to larger $Q$-ensembles has been shown to be beneficial for further control of overestimation (Chen et al., 2021).
-An et al.
-(2021) similarly apply taking a minimum over a $Q$-ensemble in model-free offline RL to mitigate the more extreme pathological value overestimation caused by the out-of-sample problem.
-Offline Model-Free Methods Model-free offline algorithms can broadly be divided into action constraint methods and value pessimism-based methods, which each represent a different approach to the out-of-sample problem.
-Action constraint methods (Kostrikov et al., 2022; Kumar et al., 2019; Fujimoto & Gu, 2021) aim to avoid using out-of-sample actions in the Bellman update by ensuring selected actions are close to the dataset behavior policy $\pi^{\beta}$.
-By contrast, value pessimism-based methods aim to directly regularize the value function such that it produces low-value estimates for out-of-sample state-actions (Kumar et al., 2020; Kostrikov et al., 2021; An et al., 2021), hence mitigating the overestimation mechanism described in Section 2.
-Of these, the most similar to our RAVL algorithm is EDAC (An et al., 2021) which applies value pessimism at any out-of-sample state-actions using minimization over a $Q$-ensemble.
-[^4] As discussed in Section 6, our RAVL approach can be viewed as reducing the required value pessimism compared to EDAC.
-Offline Model-Based Methods As alluded to in Section 1, the prevailing (incorrect) view in the offline model-based literature is that the ability to collect additional data in the model effectively resolves the out-of-sample issue, as the agent is now able to try out any previously out-of-sample behavior.
-The primary concern stated for model-based methods is therefore to account for errors in the learned dynamics model to avoid model exploitation.
-A broad class of methods explicitly penalizes the reward by the estimated level of model uncertainty (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023; Lu et al., 2022; Ball et al., 2021), typically using variance over a dynamics ensemble.
-Rigter et al.
-(2022) aim to avoid model exploitation by setting up an adversarial two-player game between the policy and model.
-Critically, all previously described methods draw their motivation from dealing with dynamics model errors.
-Finally, most related to our method, COMBO (Yu et al., 2021), penalizes value estimates for state-actions outside model rollouts.
-However, similarly to Yu et al.
-(2020), COMBO is theoretically motivated by the assumption of infinite horizon model rollouts, which we later show overlooks serious implications.
-4 The Edge-of-Reach Problem
-In the following section, we describe the central insight of our paper: the edge-of-reach problem.
-We begin by presenting the surprising result which exposes an error in current understanding.
-Next, we present our hypothesis, including intuition, empirical evidence on the main D4RL benchmark (Fu et al., 2020), and theoretical proof of its effect on offline model-based training.
-4.1 Surprising Result: Failure with the True Dynamics
-Table 1: Surprising result: Oracle denotes replacing the learned dynamics in any dynamics-penalized method (MOPO, MOReL, MOBILE) with the true dynamics (since these collapse to the same procedure with no model error).5 This change surprisingly results in failure on the D4RL MuJoCo v2 datasets.
-For comparison with using Learned dynamics (with dynamics uncertainty penalization) we list MOPO’s results (as it is the weakest of the dynamics-penalized methods).
-Statistical significance highlighted (6 seeds).
-Environment Learned Oracle
-Halfcheetah random 38.5 35.2$\pm$3.4 $\downarrow$9%
-medium 73.0 72.2$\pm$4.1 -0%
-mixed 72.1 72.2$\pm$4.1 -0%
-medexp 90.8 84.9$\pm$17.6 $\downarrow$7%
-Hopper random 31.7 9.5$\pm$1.1 $\downarrow$70%
-medium 62.8 $1$downarrow73%
-mixed 103.5 71.5$\pm$32.2 $\downarrow$31%
-medexp 81.6 6.1$\pm$6.8 $\downarrow$93%
-Walker2d random 7.4 -0.4$\pm$0.2 $\downarrow$105%
-medium 84.1 7.2$\pm$1.6 $\downarrow$91%
-mixed 85.6 7.9$\pm$1.9 $\downarrow$91%
-medexp 112.9 7.7$\pm$2.1 $\downarrow$93%
-As described in Section 1, the current understanding in offline model-based RL is that, by adding a model to enable the collection of new data throughout training, the offline problem is effectively transformed into online learning.
-With this view, the key remaining difference compared to online RL is that: in online RL, data is collected with the true dynamics, while in offline model-based RL, data is collected in an approximate dynamics model.
-Existing model-based algorithms use this understanding, and as a result base their approaches on trying to mitigate the effect of dynamics model errors, for example by introducing reward or termination penalties based on estimates of dynamics uncertainty (Yu et al., 2020; Sun et al., 2023; Kidambi et al., 2020).
-[…]
-Figure 2: An illustration of how states that are sampled only in the final step of rollouts (i.e.
-as $s_{k}$) will appear only as nextstates in the model-based training distribution.
-These states will be used as targets in the Bellman update (see Equation 1), but will not appear as $s$ in the dataset, and hence never themselves be updated.
-This mismatch leads to a breakdown of Bellman-based learning.
-4.2 The Edge-Of-Reach Hypothesis
-Our observation in Section 4.1 reveals that issues in offline model-based RL cannot be explained by dynamics model errors - a widespread misconception.
-We therefore require a different understanding.
-We begin with the main intuition: in offline model-based RL, synthetic data $\mathcal{D}_{\textrm{rollouts}}$ is generated as short $k$-step rollouts starting from states in the original dataset $\mathcal{D}_{\textrm{offline}}$ (see Algorithm 1).
-Crucially, under this procedure, there can exist some edge-of-reach states which can be reached in the final rollout step, but which cannot - under any policy - be reached earlier (see Figure 1 red).
-Sampling these states triggers a breakdown in learning, since even with the ability to collect unlimited additional data, the agent is never able to reach them ‘in time’ to try actions from them and hence never receives feedback.
-More concretely: Figure 2 illustrates how, due to being sampled only at the final step in a rollout, edge-of-reach states will be used for targets in Bellman-based learning but are never themselves updated.
-Thus, the out-of-sample action problem in model-free methods persists: since values at edge-of-reach states can be arbitrarily misestimated, incorrect values are then propagated over the entire state-action space.
-Moreover, when used with function approximation, the $\max$ operation in policy improvement can then exploit this misestimation, picking out the most heavily overestimated values (Levine et al., 2020).
-This results in extreme value overestimation, often with $Q$-values increasing unboundedly over training (as we see in Figure 3).
-Edge-of-Reach Hypothesis: The procedure of limited horizon rollouts from a fixed dataset leads to the existence of edge-of-reach states at which the agent is unable to collect corrective feedback.
-This lack of corrective feedback triggers pathological value overestimation.
-Do we expect this to actually happen in practice?
-Failure via this mode requires (a) existence of edge-of-reach states, and (b) such edge-of-reach states to be sampled.
-The typical combination of $k<<H$ along with a limited pool of starting states ($s\in\mathcal{D}_{\textrm{offline}}$) means the rollout distribution is unlikely to sufficiently cover the full state space $\mathcal{S}$, making (a) likely.
-Moreover, we observe pathological ‘edge-of-reach seeking’ behavior in which the agent appears to ‘seek out’ edge-of-reach states due to this being the source of overestimation (see Section 5), thus making (b) likely.
-In Appendix A, we give a more thorough and unified view of model-free and model-based offline RL, dividing the problem into independent conditions for states and actions.
-4.3 Definitions and Formalization
-We now formalize the intuitions presented in Section 4.2, and present theoretical grounding for the edge-of-reach problem and how errors can propagate to all transitions.
-Definition 1 (Edge-of-reach states).
-Consider a deterministic transition model $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, rollout length $k$, and some distribution over starting states $\nu_{0}$.
-For some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$, rollouts are then generated according to $s_{0}\sim\nu_{0}(\cdot)$, $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim T(\cdot|s_{t},a_{t})$ for $t=0\dots k-1$, giving $(s_{0},a_{0},s_{1},\dots,s_{k})$.
-Let us use $\rho_{t,\pi}(s)$ to denote the marginal distributions over $s_{t}$.
-We define a state $s\in S$ edge-of-reach with respect to ($T$, $k$, $\nu_{0}$) if: for $t=k$, $\exists\>\pi$ s.t. $\rho_{t,\pi}(s)>0$, but, for $t=1,\dots,k-1$ and $\forall\>\pi$, $\rho_{t,\pi}(s)=0$.
-In our case, $\nu_{0}$ is the distribution of states in $\mathcal{D}_{\textrm{offline}}$.
-With stochastic transition models (e.g.
-Gaussian models), we may have the case that no state will truly have zero density, and we relax this definition slightly to $\rho_{t,\pi}(s)<\epsilon$ for some small $\epsilon$.
-The next proposition quantifies how errors in edge-of-reach states are propagated to preceding states.
-Proposition 1 (Error propagation from edge-of-reach states).
-Consider a rollout of length $k$, $(s_{0},a_{0},s_{1},\dots,s_{k})$.
-Suppose that the state $s_{k}$ is edge-of-reach and the approximate value function $Q^{j}(s_{k},\pi(s_{k}))$ has error $\epsilon$.
-Then, standard value iteration will compound error $\gamma^{k-t}\epsilon$ to the estimates of $Q^{j+1}(s_{t},a_{t})$ for $t=1,\dots,k-1$.
-(Proof in Appendix D.)
-In practice, model rollouts are sampled in minibatches and this effect will occur on average throughout training.
-This is an analogous statement to that of Kumar et al.
-(2019) in the offline model-free setting; however, its significance in the context of model-based methods has previously been overlooked.
-In the next section, we provide further evidence of this edge-of-reach problem being the predominant issue in offline model-based RL.
-Table 2: Per-step rewards with the base offline model-based procedure (MBPO, see Algorithm 1).
-Rewards in model rollouts are close to those with the true dynamics, showing that model exploitation could not explain the subsequent value overestimation.
-Further, it indicates (as in Lu et al.
-(2022); Janner et al.
-(2019)) that the model is largely accurate for short rollouts and hence unlikely to be vulnerable to model exploitation.
-Environment Model reward True reward
-Hc-med 4.89$\pm$1.27 5.01$\pm$1.09
-Hop-mixed 2.37$\pm$1.05 2.44$\pm$1.02
-W2d-medexp 3.96$\pm$1.41 3.96$\pm$1.41
-Figure 3: The base MBPO procedure results in poor performance (left) with an exponential increase in $Q$-values (right) on the D4RL benchmark.
-Mean and standard deviation given over 6 seeds.
-Approx $Q^{*}$ roughly indicates the $Q$-value for a normalized score of 100 (with $\gamma=0.99$).
-Results are shown for Walker2d-medexp, but we note similar trends across other D4RL datasets.
-Figure 4: Experiments on the simple environment, illustrating the edge-of-reach problem and potential solutions.
-(a) Reward function, (b) final (failed) policy with naïve application of SAC, (c) final (successful) policy with patching in oracle $Q$-values for edge-of-reach states, (d) final (successful) policy with RAVL, (e) returns evaluated over training, (f) mean $Q$-values evaluated over training.
-4.4 Evidence for the Edge-of-Reach Problem
-There are two potential issues in offline model-based RL: (1) dynamics model errors and subsequent model exploitation, and (2) the edge-of-reach problem and subsequent pathological value overestimation.
-We ask: Which is the true source of the issues observed in practice?
-While (1) is stated as the sole issue and hence the motivation in prior methods, the results in Table 1 are strongly at odds with this explanation.
-Furthermore, in Table 2 we check and find that the difference between the rewards predicted by the model vs the true rewards is insignificant, which means that (1) fails to explain the value explosion in Figure 3.
-Instead, edge-of-reach-induced pathological overestimation mechanism of (2) exactly predicts this value explosion, with the observations very closely resembling those of the analogous model-free out-of-sample problem (Kumar et al., 2019).
-Furthermore, (2) is consistent with the ‘surprising result’ observations in Table 1.
-5 Analysis with a Simple Environment
-In the previous section, we provided a plausible theoretical explanation and intuition for unexpected failure observed in Section 4.1.
-In this section, we construct a simple environment and use careful analysis to empirically confirm this hypothesis.
-We first show that naïve application of actor-critic algorithms results in exploding $Q$-values and failure to learn despite using the true dynamics model.
-Next, we verify that edge-of-reach states are the source of this problem by showing that correcting value estimates only at these states is sufficient to resolve the problem.
-5.1 Setup
-[…]
-The only difference here compared to online RL is that we use a data collection horizon of $k=10$ such that $k<H$.
-This small change results in the existence of edge-of-reach states (those between the red and orange boxes in Figure 4), and hence captures the problematic component of the MBPO procedure (see Algorithm 1).
-The procedure is therefore training SAC in the true environment with limited length trajectories, exactly as is done in Experiment 4.1.
-Note: There is no notion of an approximate dynamics model, as we are demonstrating that the issues are not caused by dynamics model errors, as is currently assumed.
-We instead demonstrate that it is the existence of edge-of-reach states (due to the use of $k<H$ as per the standard MBPO procedure) that triggers pathological value overestimation.
-5.2 Observing Pathological Value Overestimation
-Since the only difference between this setup and online RL is the use of trajectory collection horizon $k<H$ (as in MBPO), we may expect SAC to learn a sensible policy.
-However, we observe that SAC fails, with $Q$-values growing unboundedly over training (see blue Figure 4(c,d)).
-Looking at the rollout trajectories sampled over training (see Figure 7) we see the following: (Before 25 epochs) Initially, the agent follows the reward function, and performance increases.
-(Between 25 and 160 epochs) Value misestimation takes over, and the policy begins to aim toward unobserved state-actions (since their values can be misestimated and hence overestimated), and performance subsequently decreases.
-(After 160 epochs) This ‘edge-of-reach seeking’ behavior compounds with each epoch, leading to the agent eventually reaching edge-of-reach states.
-From this point onwards, the agent samples edge-of-reach states at which it never receives any corrective feedback.
-The consequent pathological value overestimation results in a complete collapse in performance.
-In Figure 4(b) we visualize the final policy and see that it completely ignores the reward function.
-Instead, it aims towards an arbitrary edge-of-reach state with the highest (heavily overestimated) $Q$-value.
-5.3 Verifying the Hypothesis Using Value Patching
-Our hypothesis is that the source of this problem is value misestimation at edge-of-reach states.
-We verify this with our SAC-OraclePatch experiments, in which we replace a subset of the value estimates in the Bellman targets with their true values.
-In yellow in Figure 4 we demonstrate that applying patching solely for edge-of-reach states is sufficient to completely solve the problem.
-This is particularly compelling as in practice only a very small proportion of states had correct values patched in (0.4% of states over training).
-Of course, patching in true values is not possible in practice, thus in the next section we develop a practical approximation to this.
-We show in Section 7.1 that our proposed method, RAVL (see green in Figure 4), has a very similar effect to that of the ideal SAC-OraclePatch intervention and completely resolves the edge-of-reach problem.
-Table 3: A comprehensive evaluation of RAVL over the standard D4RL MuJoCo benchmark.
-We show the mean and standard deviation of the final performance averaged over 6 seeds.
-Our simple approach matches the state-of-the-art without any explicit dynamics penalization.
-Model-Free Model-Based
-Environment BC CQL EDAC MOPO COMBO RAMBO MOBILE RAVL (Ours)
-Halfcheetah medium 43.2 46.9 65.9 73.0 54.2 77.9 74.6 78.7$\pm$2.0
-mixed 37.6 45.3 61.3 72.1 55.1 68.7 71.7 74.9$\pm$2.0
-medexp 44.0 95.0 106.3 90.8 90.0 95.4 108.2 102.1$\pm$8.0
-Hopper medium 54.1 61.9 101.6 62.8 97.2 87.0 106.6 90.6$\pm$11.9
-mixed 16.6 86.3 101.0 103.5 89.5 99.5 103.9 103.1$\pm$1.1
-medexp 53.9 96.9 110.7 81.6 111.1 88.2 112.6 110.1$\pm$1.8
-Walker2d medium 70.9 79.5 92.5 84.1 81.9 84.9 87.7 86.3$\pm$1.6
-mixed 20.3 76.8 87.1 85.6 56.0 89.2 89.9 83.0$\pm$2.5
-medexp 90.1 109.1 114.7 112.9 103.3 56.7 115.2 115.5$\pm$2.4
-6 RAVL: Reach-Aware Value Learning
-As verified in Section 5.3, the edge-of-reach problem stems from value overestimation at edge-of-reach states.
-To resolve this issue, we therefore need to (A) detect and (B) prevent value overestimation at edge-of-reach states.
-(A) Detecting edge-of-reach states: As illustrated in Figure 2, edge-of-reach states are those that remain outside the distribution of states being updated, i.e.
-the outside training distribution of the $Q$-function, $\mathcal{D}_{\textrm{rollouts}}$.
-A natural solution to detecting out-of-distribution (OOD) inputs is to use a deep ensemble (Lakshminarayanan et al., 2017).
-As such, high variance over an ensemble of $Q$-functions can be used to detect edge-of-reach states.
-We demonstrate that this is effective in Figure 5.
-Note this is different to related model-free methods since the distribution being considered is that of $\mathcal{D}_{\textrm{rollouts}}$ rather than $\mathcal{D}_{\textrm{offline}}$.
-(B) Preventing overestimation at these states: Once we have detected edge-of-reach states, we may simply apply value pessimism methods from the offline model-free literature to these states.
-Our choice of an ensemble for part (A) conveniently allows us to simply minimize over an ensemble of $Q$-functions, which effectively penalizes values based on ensemble variance.
-Our resulting proposal is Reach-Aware Value Learning (RAVL).
-Concretely, RAVL simply involves training an ensemble of $N_{\textrm{critic}}$ $Q$-functions similarly to EDAC (An et al., 2021), indexed by $n$, and modifying Equation 1 to:
-$\small\begin{split}Q_{n}^{j+1}\leftarrow\underset{Q_{n}}{\arg\min}\;\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D},a^{\prime}\sim\pi^{j}(\cdot\mid s^{\prime})}[\qquad\qquad\>\\ (\underbrace{Q_{n}(s,a)}_{\text{\tiny input}}-\underbrace{[r+\gamma\min_{n=1,\dots,N_{\textrm{critic}}}Q_{n}^{j}(s^{\prime},a^{\prime})]}_{\text{\tiny Bellmantarget}})^{2}\>]\end{split}$ (2)
-As is standard in offline model-based RL, we optimize this objective using MBPO (Janner et al., 2019) and SAC (Haarnoja et al., 2018) as in Algorithm 1.
-RAVL is therefore a solution to offline model-based RL, which relies solely on value pessimism ideas from the model-free literature.
-This is in stark contrast to typical methods which use dynamics uncertainty penalization and therefore also provides further evidence that the edge-of-reach problem is the true source of issues.
-RAVL represents combining model-based training with the current SOTA model-free algorithm EDAC.
-The addition of the model compared to EDAC can be seen as relaxing the value pessimism requirements from any state-actions OOD with respect to $\mathcal{D}_{\textrm{offline}}$, to those OOD with respect to $\mathcal{D}_{\textrm{rollouts}}$.
-This perspective highlights our algorithm’s simplicity, with the approach being made possible due to our corrected understanding of the true issues involved in model-based offline RL.
-Our subsequently unified view of model-based and model-free approaches has allowed us to transfer ideas from the model-free literature for dealing with the out-of-sample problem and directly apply them to solve the edge-of-reach issue in the model-based setting.
-7 Empirical Evaluation
-In this section, we first evaluate RAVL on the simple environment from Section 5.
-Next, we show that RAVL scales, achieving strong performance on the standard D4RL benchmark and a new state-of-the-art on challenging pixel-based V-D4RL tasks.
-Finally, we compare RAVL to prior dynamics uncertainty penalized methods and elucidate why they may work despite not explicitly addressing the edge-of-reach problem.
-We provide full hyperparameters and a discussion of run-time in Section C.2.
-7.1 Evaluation on the Simple Environment
-We first test RAVL on the simple environment from Section 5 (see green in Figure 4).
-We observe that it achieves the same result as the theoretically optimal but practically impossible method of patching in the true values at edge-of-reach states, SAC-OraclePatch.
-The $Q$-values are stabilized and match those of SAC-OraclePatch.
-Moreover, in Figure 5, we demonstrate that RAVL can precisely detect edge-of-reach states as they have significantly higher ensemble variance than within-reach states, as expected.
-Figure 5: RAVL’s effective penalty of variance over the $Q$-value ensemble on the environment in Section 5.
-As expected, values at edge-of-reach states have significantly higher variance than for within-reach states, meaning, exactly as desired, RAVL is able to detect and apply a value pessimism penalty at edge-of-reach states.
-7.2 Evaluation on the D4RL Benchmark
-Next, we show that RAVL scales to the standard offline RL benchmark, D4RL (Fu et al., 2020).
-We consider the MuJoCo (Todorov et al., 2012) v2 datasets.
-Figure 3 shows RAVL successfully stabilizes training, and Table 3 shows that we can match the performance of the current top model-based method, MOBILE (Sun et al., 2023), without the need for any explicit dynamics penalization.
-This is beneficial in terms of runtime, as MOBILE requires multiple transition samples to calculate its penalty.
-Compared to EDAC (An et al., 2021), the model-free equivalent of our algorithm, RAVL achieves much higher and state-of-the-art performance on the Halfcheetah mixed and medium datasets.
-This demonstrates the clear benefit of generating additional model-based samples.
-Our approach is far stronger than COMBO (Yu et al., 2021) which introduced a less effective form of value conservatism into model-based methods.
-7.3 Evaluation on Pixel-Based tasks
-The insights from RAVL also lead to improvements in performance on the challenging pixel-based benchmark V-D4RL (Lu et al., 2023).
-This uses the Offline DreamerV2 algorithm which trains on trajectories in a learned latent space, thus demonstrating that the edge-of-reach problem is more general and widespread.
-We observe in Table 4 that RAVL gives a strong boost in performance (+20.3% overall) on the medium and medexp level datasets, while not helping in the more diverse random and mixed datasets.
-This observation fits with our intuition of the edge-of-reach problem: the medium and medexp level datasets are likely to have less coverage of the state space, and thus we would expect them to suffer more from edge-of-reach issues and the ‘edge-of-reach seeking behavior’ demonstrated in Section 5.2.
-We note also that these results are without the ensemble diversity regularizer from EDAC (An et al., 2021) (used on the D4RL benchmark), which we anticipate may further increase performance.
-Table 4: We show that RAVL extends to the challenging pixel-based V-D4RL benchmark. The change in mean performance is $+20.3\%$. Mean and standard deviation given over 6 seeds.
-Environment Offline DV2 RAVL (Ours)
-Walker-walk medium 34.1$\pm$19.7 48.6$\pm$8.7
-medexp 43.9$\pm$34.4 47.6$\pm$18.3
-Cheetah-run medium 17.2$\pm$3.5 18.3$\pm$4.3
-medexp 10.4$\pm$3.5 12.6$\pm$4.4
-7.4 Reinterpretation of Prior Methods
-Prior model-based methods were unaware of the edge-of-reach problem, instead being motivated solely by mitigating issues due to model errors.
-However, our investigation strongly suggests that the edge-of-reach problem is likely to be the true main underlying issue on the D4RL benchmark (see Section 4.4).
-This leads us to ask: What allows prior model error-motivated methods to work despite ignoring the critical edge-of-reach problem?
-We find that, while ostensibly to address model errors, existing dynamics penalties may accidentally help to address the edge-of-reach problem, as we observe a positive correlation between the penalties used in dynamics uncertainty methods and RAVL’s effective penalty of value ensemble variance (see Figure 6 in Section B.1).
-This may be expected, as dynamics uncertainty will naturally be higher further away from $\mathcal{D}_{\textrm{offline}}$, which is also where edge-of-reach states are more likely to lie.
-We additionally note that, since the edge-of-reach problem is orthogonal to dynamics misestimation, RAVL could be combined with appropriate model uncertainty penalization (Yu et al., 2020; Sun et al., 2023; Lu et al., 2022) for environments with less accurate models.
-8 Conclusion
-In this paper, we expose a major misunderstanding in current model-based offline RL.
-We provide strong evidence that - contrary to common understanding - the true source of issues experienced by current methods is the edge-of-reach problem, a previously unnoticed issue whereby the standard offline model-based RL procedure leads to the existence of a set of edge-of-reach states that triggers pathological value misestimation and collapse of learning.
-Exposing this reveals a strong connection between model-free and model-based offline RL, which allows us to unify our understanding of these two areas.
-Based on this new unified perspective, we are able to transfer ideas from model-free literature and directly apply them to solve the edge-of-reach problem in model-based methods.
-Our resulting practical algorithm RAVL, is simple, robust, and achieves strong performance across both proprioceptive and pixel-based benchmarks.
-We open-source our code and hope that the insights presented here may help further ideas to be transferred between model-free and model-based methods.
-Finally, a more complete theoretical picture could be obtained by transferring finite horizon rollout theory from online model-based reinforcement learning to the offline setting (Janner et al., 2019).
-Statement of Broader Impact
-This paper presents work whose goal is to advance the field of offline reinforcement learning.
-Since we evaluate on simulated control problems in this paper, we do not expect any specific ethical concerns to arise.
-However, when offline reinforcement learning is deployed in the real world, it is important to follow responsible ethical principles in ensuring the trustworthiness and reliability of our systems.
-Acknowledgements
-Anya Sims and Cong Lu are funded by the Engineering and Physical Sciences Research Council (EPSRC).
-The authors would like to thank the anonymous Agent Learning in Open-Endedness Workshop at NeurIPS 2023 and ICLR 2024 reviewers for positive and constructive feedback which helped to improve the paper.
-We would also like to thank Philip J.
-Ball and Shimon Whiteson for helpful feedback and discussions on earlier drafts of this work.
-Supplementary Material
 Appendix A A Unified Perspective of Model-Based and Model-Free RL
@@
 We abbreviate $x$ is in distribution with respect to $\rho$ as $x\in^{\textrm{dist}}\rho$.
-A.2 $Q$-Learning Conditions
+A.2 $\mathbf{Q}$-Learning Conditions
 As described in Section 2, given some policy $\pi$, we can attempt to learn the corresponding $Q$-function with the following iterative process:
@@
 A.3 Comparison Between Model-Free and Model-Based Methods
-Table 5: A summary of the comparison between model-free and model-based offline RL in relation to the conditions on $Q$-learning.
-Action condition violation State condition violation Main source of issues
-Model-free Possible, common in practice Possible, but uncommon in practice $\rightarrow$ Action condition violation
-Model-based Not possible Possible, common in practice $\rightarrow$ State condition violation
 In offline model-free RL, $\mathcal{D}=\mathcal{D}_{\textrm{offline}}$, with $\pi^{\mathcal{D}}=\pi^{\beta}$.
@@
 By contrast, the mismatch between $\pi$ (used to sample $a^{\prime}$ in $Q$-learning) and $\pi^{\beta}$ (used to sample $a$ in the dataset $\mathcal{D}_{\textrm{offline}}$) often does lead to significant violation of the action condition.
-This exacerbates the overestimation bias in $Q$-learning (see Section 3), and can result in pathological training dynamics and $Q$-value explosion over training (Kumar et al., 2019).
+This exacerbates the overestimation bias in $Q$-learning (see Section 7), and can result in pathological training dynamics and $Q$-value explosion over training [18].
 On the other hand, in offline model-based RL, the dataset $\mathcal{D}=\mathcal{D}_{\textrm{rollouts}}$ is collected on-policy according to the current (or recent) policy such that $\pi^{\mathcal{D}}\approx\pi$.
@@
 Our paper then considers the pathological case of edge-of-reach states, which will always violate the state condition.
-Appendix B Additional Figures
-B.1 Comparison to Prior Approaches
-In Figure 6, we plot the dynamics uncertainty-based penalty used in MOPO (Yu et al., 2020) against the effective penalty in RAVL of variance of the value ensemble.
-The positive correlation between MOPO’s penalty and the edge-of-reach states targeting penalty of RAVL helps to explain why prior methods may work despite not considering the crucial edge-of-reach problem.
-Figure 6: We find that the dynamics uncertainty-based penalty used in MOPO (Yu et al., 2020) is positively correlated with the variance of the value ensemble of RAVL, suggesting prior methods may unintentionally address the edge-of-reach problem.
-Pearson correlation coefficients are 0.49, 0.43, and 0.27 for Hopper-mixed, Walker2d-medexp, and Halfcheetah-medium respectively.
-B.2 Visualizations of Rollouts Through Training
-In Figure 7, we provide a visualization of the rollouts sampled over training in the simple environment for each of the algorithms analyzed in Figure 4 (see Section 5).
-This accompanies the discussion of the behavior of SAC during training in Section 5.2.
-Figure 7: A visualization of the rollouts sampled over training on the simple environment in Section 5.
-We note the pathological behavior of the baseline SAC, and the success of the ideal intervention SAC-OraclePatch, and our practically realizable method RAVL.
+This comparison between model-based and model-free is summarized in LABEL:tab:conditions1
+Table 3: A summary of the comparison between model-free and model-based offline RL in relation to the conditions on $Q$-learning as described in Appendix A.
+Action condition violation State condition violation Main source of issues
+Model-free Yes, common in practice Yes, but uncommon in practice $\rightarrow$ Action condition violation
+Model-based No Yes, common in practice $\rightarrow$ State condition violation
+Appendix B Error Propagation Result
+Proposition 1 (Error propagation from edge-of-reach states).
+Consider a rollout of length $k$, $(s_{0},a_{0},s_{1},\dots,s_{k})$.
+Suppose that the state $s_{k}$ is edge-of-reach and the approximate value function $Q^{j}(s_{k},\pi(s_{k}))$ has error $\epsilon$.
+Then, standard value iteration will compound error $\gamma^{k-t}\epsilon$ to the estimates of $Q^{j+1}(s_{t},a_{t})$ for $t=1,\dots,k-1$.
+(Proof in Appendix B.)
+B.1 Proof
+In this section, we provide a proof of Proposition 1. Our proof follows analogous logic to the error propagation result of Kumar et al. [18].
+Proof.
+Let us denote $Q^{\ast}$ as the optimal value function, $\zeta_{j}(s,a)=|Q_{j}(s,a)-Q^{\ast}(s,a)|$ the error at iteration $j$ of Q-Learning, and $\delta_{j}(s,a)=|Q_{j}(s,a)-\mathcal{T}Q_{j-1}(s,a)|$ the current Bellman error.
+Then first considering the $t=k-1$ case,
+$\zeta_{j}(s_{t},a_{t})$ $=|Q_{j}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (4)
+$=|Q_{j}(s_{t},a_{t})-\mathcal{T}Q_{j-1}(s_{t},a_{t})+\mathcal{T}Q_{j-1}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (5)
+$\leq|Q_{j}(s_{t},a_{t})-\mathcal{T}Q_{j-1}(s_{t},a_{t})|+|\mathcal{T}Q_{j-1}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (6)
+$=\delta_{j}(s_{t},a_{t})+\gamma\zeta_{j-1}(s_{t+1},a_{t+1})$ (7)
+$=\delta_{j}(s_{t},a_{t})+\gamma\epsilon$ (8)
+Thus the errors at edge-of-reach states are discounted and then compounded with new errors at $Q^{j}(s_{k-1},a_{k-1})$.
+For $t<k-1$, the result follows from repeated application of Equation 7 along the rollout.
+B.2 Extension to Stochastic Environments and Practical Implementation
+With stochastic transition models (e.g.
+Gaussian models), we may have the case that no state will truly have zero density, in which case we relax the definition of edge-of-reach states slightly (see Section 3.3) from $\rho_{t,\pi}(s)=0$ to $\rho_{t,\pi}(s)<\epsilon$ for some small $\epsilon$.
+During optimization, in practice, model rollouts are sampled in minibatches and thus the above error propagation effect will occur on average throughout training.
+An analogous model-free statement has been given in Kumar et al.
+[18]; however, its significance in the context of model-based methods was previously not considered.
 Appendix C Implementation Details
-In this section, we give full implementation details for RAVL.
+In this section, we provide full implementation details for RAVL.
 C.1 Algorithm
-As is typical in offline model-based reinforcement learning (Yu et al., 2020; Lu et al., 2022; Kidambi et al., 2020; Sun et al., 2023), we use short MBPO (Janner et al., 2019) style rollouts.
-The agent is trained using the EDAC (An et al., 2021) losses, a conservative ensemble-based algorithm based on SAC (Haarnoja et al., 2018).
-Our transition model follows the standard setup in model-based offline algorithms, being realized as a deep ensemble (Chua et al., 2018) and trained via maximum likelihood estimation.
-This is summarized in the pseudocode for offline model-based methods given in Algorithm 1.
+We use the base model-based procedure as given in Algorithm 1 and shared across model-based offline RL methods [36, 24, 14, 29].
+This involves using short MBPO-style [13] model rollouts to train an agent based on SAC [11].
+We modify the SAC agent with the value pessimism losses of EDAC [1].
+Our dynamics model follows the standard setup in model-based offline algorithms, being realized as a deep ensemble [5] and trained via maximum likelihood estimation.
 C.2 Hyperparameters
-For the D4RL (Fu et al., 2020) MuJoCo datasets presented in Section 7.2, we sweep over the following hyperparameters and list the optimal choices in Table 6.
+For the D4RL [7] MuJoCo results presented in Table 2, we sweep over the following hyperparameters and list the choices used in Table 4.
+“Base” refers to the shared base procedure in model-based offline RL shown in Algorithm 1.
   • (EDAC) Number of $Q$-ensemble elements $N_{\textrm{critic}}$, in the range $\{10,50\}$
   • (EDAC) Ensemble diversity weight $\eta$, in the range $\{1,10,100\}$
-  • (MBPO) Model rollout length $k$, in the range $\{1,5\}$
-  • (MBPO) Real-to-synthetic data ratio $r$, in the range $\{0.05,0.5\}$
+  • (Base) Model rollout length $k$, in the range $\{1,5\}$
+  • (Base) Real-to-synthetic data ratio $r$, in the range $\{0.05,0.5\}$
 The remaining model-based and agent hyperparameters are given in Table 5.
-Almost all environments use a small $N_{\textrm{critic}}=10$ similar to An et al.
-(2021), with only the Hopper-medexp dataset needing $N_{\textrm{critic}}=30$.
+Almost all environments use a small $N_{\textrm{critic}}=10$, with only the Hopper-medexp dataset needing $N_{\textrm{critic}}=30$.
+For the uncertainty-free dynamics model experiments in Table 1 we take the same hyperparameters as for the learned dynamics model for $N$, $k$, and $r$, and try different settings for $\eta\in\{1,10,100,200\}$.
+The choices used are shown in Table 4.
+Note that for existing methods based on dynamics penalties, the analogous hyperparameter (penalty weighting) has no effect under the true dynamics model as the penalties will always be zero (since dynamics uncertainty is zero).
+For the experiments in Figure 1 with intermediate dynamics model accuracies, we tune the ensemble diversity coefficient $\eta$ over $\eta\in\{1,10,100,200\}$ for RAVL, and analogously we tune the uncertainty penalty coefficient $\lambda$ over $\lambda\in\{1,5,10,100,200\}$ for MOPO.
 Our implementation is based on the Clean Offline Reinforcement Learning (CORL, Tarasov et al.
-(2022)) repository, released at https://github.com/tinkoff-ai/CORL under an Apache-2.0 license.
-Our algorithm takes approximately 1 day to run on the standard D4RL datasets using an A100 GPU for the full number of epochs.
-Table 6: Variable hyperparameters for RAVL used in D4RL MuJoCo locomotion tasks.
+[33]) repository, released at https://github.com/tinkoff-ai/CORL under an Apache-2.0 license.
+Our algorithm takes on average 6 hours to run using a V100 GPU for the full number of epochs.
+Table 4: Variable hyperparameters for RAVL used in D4RL MuJoCo locomotion tasks with the learned dynamics (see Table 2).
+With the uncertainty-free dynamics (see Table 1), we tune only $\eta$, with the settings used shown in brackets.
       Environment       $\mathbf{N_{critic}}$       $\bm{\eta}$       $\mathbf{k}$       $\mathbf{r}$
-HalfCheetah medium 10 1 5 0.05
-mixed 10 100 5 0.05
-medexp 10 1 5 0.5
-Hopper medium 10 100 1 0.5
-mixed 10 10 1 0.5
-medexp 50 100 1 0.5
-Walker2d medium 10 10 1 0.5
-mixed 10 10 1 0.5
-medexp 10 1 1 0.5
+      HalfCheetah       random       10       10 (1)       5       0.5
+      medium       10       1 (1)       5       0.05
+      mixed       10       100 (1)       5       0.05
+      medexp       10       1 (1)       5       0.5
+      Hopper       random       10       100 (1)       5       0.05
+      medium       10       100 (10)       1       0.5
+      mixed       10       10 (10)       1       0.5
+      medexp       30       100 (10)       1       0.5
+      Walker2d       random       10       1 (100)       5       0.05
+      medium       10       10 (100)       1       0.5
+      mixed       10       10 (100)       1       0.5
+      medexp       10       1 (200)       1       0.5
 Table 5: Fixed hyperparameters for RAVL used in D4RL MuJoCo locomotion tasks.
@@
 C.3 Pixel-Based Hyperparameters
-For the V-D4RL (Lu et al., 2023) DeepMind Control Suite datasets presented in Section 7.3, we use the default hyperparameters for the Offline DV2 algorithm, which are given in Table 8.
+For the V-D4RL [25] DeepMind Control Suite datasets presented in Table 9, we use the default hyperparameters for the Offline DV2 algorithm, which are given in Table 6.
 We found keeping the uncertainty weight $\lambda=10$ improved performance over $\lambda=0$ which shows RAVL can be combined with dynamics penalty-based methods.
@@
         number of critics ($\textrm{N}_{\textrm{critic}}$)         5
-We swept over $\{2,5,20\}$ for $\textrm{N}_{\textrm{critic}}$ but found a single value of $5$ worked well for all environments we consider.
-Appendix D Proof of Error Propagation Result
-In this section, we provide a proof of Proposition 1. Our proof follows analogous logic to the error propagation result of Kumar et al. (2019).
-Proof.
-Let us denote $Q^{\ast}$ as the optimal value function, $\zeta_{j}(s,a)=|Q_{j}(s,a)-Q^{\ast}(s,a)|$ the error at iteration $j$ of Q-Learning, and $\delta_{j}(s,a)=|Q_{j}(s,a)-\mathcal{T}Q_{j-1}(s,a)|$ the current Bellman error.
-Then first considering the $t=k-1$ case,
-$\zeta_{j}(s_{t},a_{t})$ $=|Q_{j}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (4)
-$=|Q_{j}(s_{t},a_{t})-\mathcal{T}Q_{j-1}(s_{t},a_{t})+\mathcal{T}Q_{j-1}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (5)
-$\leq|Q_{j}(s_{t},a_{t})-\mathcal{T}Q_{j-1}(s_{t},a_{t})|+|\mathcal{T}Q_{j-1}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (6)
-$=\delta_{j}(s_{t},a_{t})+\gamma\zeta_{j-1}(s_{t+1},a_{t+1})$ (7)
-$=\delta_{j}(s_{t},a_{t})+\gamma\epsilon$ (8)
-Thus the errors at edge-of-reach states are discounted and then compounded with new errors at $Q^{j}(s_{k-1},a_{k-1})$.
-For $t<k-1$, the result follows from repeated application of Equation 7 along the rollout.
-Appendix E Summary of Setups Used in Comparisons
+We used a hyperparameter sweep over $\{2,5,20\}$ for $\textrm{N}_{\textrm{critic}}$ but found a single value of $5$ worked well for all environments we consider.
+Appendix D Additional Tables
+D.1 Existing Methods’ Penalties All Collapse to Zero Under the True Dynamics
+For completeness, in Table 7 we include the expressions for the various dynamics penalties proposed by the most popular model-based approaches.
+These penalties are all based on estimates of dynamics model uncertainty.
+Since there is no uncertainty under the true model, all these penalties collapse to zero with the true dynamics.
+Table 7: We show here how all existing dynamics penalized offline MBRL algorithms reduce to the same base procedure when estimated (epistemic) model uncertainty is zero (as it is with the true dynamics).
+Offline MBRL Algorithm Penalty Penalty with the true dynamics
+MOPO [36] ${\max}_{i=1,\dots,N}||\Sigma^{i}_{\phi}(s,a)||_{\mathrm{F}}$ 0
+MOReL [14] ${\max}_{1\leq i,j\leq N}||\mu^{i}_{\phi}(s,a)-\mu^{j}_{\phi}(s,a)||_{2}$ 0
+MOBILE [29] $\textrm{Std}_{i=1,\dots,N}\{\hat{T}^{i}(s,a)\}$ 0
+D.2 Model Errors Cannot Explain the $Q$-value Overestimation
+In Figure 3 we observe that the base offline model-based procedure results in $Q$-values exploding unboundedly to beyond $10^{10}$ over training.
+Existing methods attribute this to dynamics model errors and the consequent exploitation of state-action space areas where the model incorrectly predicts high reward.
+In Table 8 we examine the rewards collected by an agent in the learned dynamics model.
+Under the ‘model errors’ explanation, we would expect to see rewards sampled by the agent in the learned model to be significantly higher than under the true environment.
+Furthermore, to explain the $Q$-values being on the order of $10^{10}$, with $\gamma=0.99$ the agent would need to sample rewards on the order of $10^{8}$.
+However, the per-step rewards sampled by the agent are on the order of $10^{0}$ (and are not higher than with the true environment).
+Model errors and model exploitation therefore cannot explain the explosion in $Q$-values seen over training.
+By contrast, the edge-of-reach problem exactly predicts this extreme value overestimation behavior.
+Table 8: Per-step rewards with the base offline model-based procedure (see Algorithm 1).
+Rewards in model rollouts are close to those with the true dynamics, showing that model exploitation could not explain the subsequent value overestimation.
+Further, it indicates (as in Lu et al.
+[24], Janner et al.
+[13]) that the model is largely accurate for short rollouts and hence is unlikely to be vulnerable to model exploitation.
+            Environment             Model reward             True reward
+            Hc-med             4.89$\pm$1.27             5.01$\pm$1.09
+            Hop-mixed             2.37$\pm$1.05             2.44$\pm$1.02
+            W2d-medexp             3.96$\pm$1.41             3.96$\pm$1.41
+Appendix E Additional Visualizations
+E.1 Comparison to Prior Approaches
+In Figure 6, we plot the dynamics uncertainty-based penalty used in MOPO [36] against the effective penalty in RAVL of variance of the value ensemble.
+The positive correlation between MOPO’s penalty and the edge-of-reach states targeting penalty of RAVL indicates why prior methods may work despite not considering the crucial edge-of-reach problem.
+Figure 6: We find that the dynamics uncertainty-based penalty used in MOPO [36] is positively correlated with the variance of the value ensemble of RAVL, suggesting prior methods may unintentionally address the edge-of-reach problem.
+Pearson correlation coefficients are 0.49, 0.43, and 0.27 for Hopper-mixed, Walker2d-medexp, and Halfcheetah-medium respectively.
+E.2 Behaviour of Rollouts Over Training
+In Figure 7, we provide a visualization of the rollouts sampled over training in the simple environment for each of the algorithms analyzed in Figure 4 (see Section 4).
+This accompanies the discussion in Section 4.2 of the behavior over training.
+Figure 7: A visualization of the rollouts sampled over training on the simple environment in Section 4.
+We note the pathological behavior of the baseline, and then the success of the ideal intervention Base-OraclePatch, and our practically realizable method RAVL.
+Appendix F Evaluation on the Pixel-Based V-D4RL Benchmark
+The insights from RAVL also lead to improvements in performance on the challenging pixel-based benchmark V-D4RL [25].
+The base procedure in this pixel-based setting uses the Offline DreamerV2 algorithm of training on trajectories in a learned latent space.
+We observe in Table 9 that RAVL gives a strong boost in performance on the medium and medexp level datasets, while helping less in the more diverse random and mixed datasets.
+This observation fits with our intuition of the edge-of-reach problem: the medium and medexp level datasets are likely to have less coverage of the state space, and thus we would expect them to suffer more from edge-of-reach issues and the ‘edge-of-reach seeking behavior’ demonstrated in Section 4.2.
+The performance improvements over Offline DreamerV2 and other model-based baselines including LOMPO [26] therefore suggests that the edge-of-reach problem is general and widespread.
+We also note that these results are without the ensemble diversity regularizer from EDAC [1] (used on the D4RL benchmark), which we anticipate may further increase performance.
+Table 9: We show that RAVL extends to the challenging pixel-based V-D4RL benchmark, suggesting the out-of-reach problem is also present in the latent-space setting used by the base procedure for pixel-based algorithms.
+Mean and standard deviation given over 6 seeds.
+      V-D4RL Environment       Offline DV2       LOMPO       RAVL (Ours)
+      Walker-walk       random       28.7$\pm$13.0       21.9$\pm$8.1       29.7$\pm$2.4
+      mixed       56.5$\pm$18.1       34.7$\pm$19.7       49.3$\pm$11.3
+      medium       34.1$\pm$19.7       43.4$\pm$11.1       48.6$\pm$8.7
+      medexp       43.9$\pm$34.4       39.2$\pm$19.5       47.6$\pm$18.3
+      Cheetah-run       random       31.7$\pm$2.7       11.4$\pm$5.1       36.9$\pm$5.0
+      mixed       61.6$\pm$1.0       36.3$\pm$13.6       57.6$\pm$1.4
+      medium       17.2$\pm$3.5       16.4$\pm$8.3       18.3$\pm$4.3
+      medexp       10.4$\pm$3.5       11.9$\pm$1.9       12.6$\pm$4.4
+      Overall       35.5$\pm$12.0       26.9$\pm$10.9       37.6$\pm$7.0
+Appendix G Runtime and Hyperparameter Sensitivity Ablations
+G.1 Runtime
+We compare RAVL to SOTA method MOBILE.
+The additional requirement for RAVL compared to MOBILE is that RAVL requires increasing the $Q$-ensemble size: from $N=2$ in MOBILE to $N=10$ in RAVL (for all environments except one where we use $N=30$).
+The additional requirement for MOBILE compared to RAVL is that MOBILE requires multiple additional forwards passes through the model for each update step (for computing their dynamics uncertainty penalty).
+In Table 10 we show that ensemble size can be scaled with very minimal effect on the runtime.
+Even increasing the ensemble size to $N=100$ (far beyond the maximum ensemble size used by RAVL) only increases the runtime by at most 1%.
+Table 7 of MOBILE reports that MOBILE’s requirement of additional forward passes increases runtime by around 14%.666 Note that the times reported for EDAC in Table 7 of MOBILE are not representative of practice, since these are with a non-vectorized $Q$-ensemble (confirmed from their open-source code), which would require just a simple change to  5 lines of code.
+Table 10: Timing experiments showing that ensemble size for standard vectorized ensembles can be scaled with extremely minimal effect on runtime.
+Runtime increases by just 1% with $Q$-ensemble size increased from $N=2$ (as used in the base procedure SAC agent) to $N=100$.
+RAVL uses $N=10$ (for all environments except one where we use $N=30$), meaning the runtime increase compared to the base model-based procedure is almost negligible.
+Other methods add more computationally expensive changes to the base procedure, hence making RAVL significantly faster than SOTA (see Section G.1).
+     Ensemble size ($N$) 2 10 50 100
+Time relative to $N=2$ (baseline) ×$\times 1.00\pm 0.03$$\times 1.01\pm 0.04$$\times 1.01\pm 0.02$
+G.2 Hyperparameter Sensitivity Ablations
+In Table 11 we include results with different settings of RAVL’s ensmble diversity regularizer hyperparameter $\eta$.
+We note a minimal change in performance with a sweep across two orders of magnitude.
+This is an indication that RAVL may be applied to new settings without much tuning.
+For RAVL’s second hyperparameter of $Q$-ensemble size $N$, we note that in the main benchmarking of RAVL against other offline methods (see Table 2) we use the same value of $N=10$ across all except one environment.
+This again indicates that RAVL is robust and should transfer to new settings with minimal hyperparameter tuning.
+Table 11: Ablations results over RAVL’s diversity regularizer hyperparameter $\eta$.
+         Hyperparameter $\eta$          Halfcheetah medium          Halfcheetah mixed
+         1.0          $78.7\pm 2.0$          $61.7\pm 5.6$
+         10.0          $74.8\pm 1.5$          $73.8\pm 1.9$
+         100.0          $72.1\pm 2.2$          $74.9\pm 2.0$
+Appendix H Summary of Setups Used in Comparisons
 Throughout the paper, we compare several different setups in order to identify the true underlying issues in model-based offline RL.
 We provide a summary of them in Table 12.
-More comprehensive descriptions of each are given in the main text and in relevant table and figure captions.
+More comprehensive descriptions of each are given in the relevant positions in the main text and table and figure captions.
 Table 12: We summarize the various setups used for comparisons throughout the paper.
@@
 All methods use $k$-step rollouts from the offline dataset (or from a fixed starting state distribution in the case of the simple environment).
-Use of the MOPO penalty refers to implementing the method in Yu et al.
-(2020).
-Dynamics Model Agent Figures and Tables
+Figures and Tables Name Used Dynamics Model Agent
 Type Penalty $N_{\textrm{critic}}$ $\eta$
-Learned Ensemble ✓(MOPO) $2$ 00 Table 1, Figure 6
-Oracle True n/a $2$ 00 Table 1
-MBPO (Offline) Ensemble $2$ 00 Tables 2, Figure 3
-RAVL (Ours) Ensemble $>2$ $>0$ Tables 2, 3, Figure 3
-SAC (Truncated Horizon)* True n/a $2$ 00 Figures 4, 7
-RAVL (Ours)* True n/a $>2$ $>0$ Figures 4, 5, 7
+Table 1, Figure 6 ‘Approximate’ Ensemble ✓(Best of Table 7) $2$ 00
+Table 1 ‘True’ True n/a $2$ 00
+Tables 8, Figure 3 ‘Base’ Ensemble $2$ 00
+Tables 8, 2, Figure 3 ‘RAVL (Ours)’ Ensemble $>2$ $>0$
+Figures 4, 7 ‘Base’* True n/a $2$ 00
+Figures 4, 5, 7 ‘RAVL (Ours)’* True n/a $>2$ $>0$